### v1

# Raman Microspectroscopy for Real-Time Structure Indicator in Ultrafast Laser Writing

Abstract
Raman microspectroscopy provides non-destructive, structure-specific contrast suited to in situ monitoring of femtosecond (fs) laser fabrication processes. In this work, we introduce a sample-agnostic Raman workflow that screens candidate spectral indicators using drift-resistant, self-normalized metrics and, where appropriate, label-free hyperspectral unmixing to extract endmembers and validate metric selection. Using diamond as an example, hyperspectral maps and electrical tests show that depletion of the 1332 cm-1 sp3 line is a monotonic, robust predictor of resistance—counterintuitively outperforming sp2-based measures often assumed optimal. This enables a practical implementation: a single narrowband channel at the selected marker provides real-time feedback without full-spectrum acquisition. Because the selection hinges on general band definitions and functional validation rather than diamond-specific physics, the approach translates Raman phase information directly into device specifications and is transferable to other Raman-active hosts for in-line quality control of laser-written structures.
I Introduction
Femtosecond laser writing enables direct, three-dimensional patterning of embedded features in transparent hosts via nonlinear absorption—multi-phonon ionization (MPI) and Zener breakdown—that initiates localized structural transformation [1, 2, 3]. These processes permit buried and surface carbon wires and electrode architectures with arbitrary geometry for electrically conductive devices [4], including radiation detectors [5]. Practical subsurface fabrication benefits from pre-compensating high-index–induced spherical aberration using adaptive optics such as a spatial light modulator (SLM) to restore a tight focal volume [6, 7]. While the resulting microstructures have been examined by destructive or ex-situ methods—TEM [8, 9], SEM [2], and X-ray diffraction microscopy (XRD) [10]—their complexity and inability to operate in situ limit process monitoring.
Raman microspectroscopy is a vibrational spectroscopic technique based on the inelastic scattering of monochromatic light. It has been extensively used for non-destructive identification, structural characterization, and monitoring of chemical and physical properties across materials science, chemistry, and biology[11]. In carbon systems in particular, Raman can distinguish bonding configurations, crystallinity, and local strain, making it suitable for non-invasive, in situ, spatially resolved monitoring during fabrication [12].
To obtain robust, comparable indicators from spectra acquired across fields of view and samples, previous studies have adopted a deconvolution and normalization strategy that favors self-referenced, phase-aware metrics over single-peak intensities [11]. Absolute intensities are sensitive to power drift, focus/NA mismatch, baseline fluorescence, and strain; by contrast, combining phase-informative windows and leveraging label-free spectral deconvolution stabilizes the readout and enables cross-instrument comparison. In practice, we explicitly assign bands, normalize to pristine-material pixels, and compute complementary indicators of conversion (ratios and depletion measures defined below), converting hyperspectral maps into quantitative metrics suitable for in situ monitoring [12].
Diamond is an instructive case for Raman-guided fs fabrication. As sp3-bonded carbon, it is an electrical insulator with exceptional hardness, high thermal conductivity, a wide bandgap, and chemical robustness, enabling applications in high-power electronics, photonics, and radiation detection. Point defects such as nitrogen vacancy centers further extend functionality to quantum information and nanoscale sensing [13, 14, 15, 16]. The broader carbon design space—sp2-bonded forms (graphite, graphene, carbon nanotubes) with delocalized electrons for high electrical/thermal conductivity [17, 18], allotropic variants such as hexagonal diamond (lonsdaleite) [19], and mixed-phase nanocomposites such as ’diaphite’ [9, 20, 21]—underscores the value of a phase-resolved optical readout. Although diamond is transparent at 520 nm (2.38 eV) due to its 5.47 eV bandgap, which prevents linear photon absorption from the writing laser, the fabrication process instead relies on nonlinear absorption involving multi-phonon ionization (MPI) and Zener breakdown, initiating a confined sp3 to sp2 phase transition [1, 2, 3], forming local graphitized regions with device-relevant geometries.
Here, we use Raman microspectroscopy to systematically assess fs-laser-written graphitic electrodes in diamond. Leveraging these vibrational modes, we perform direct mapping of surface graphitization through explicit band assignment and label-free spectral deconvolution, and we correlate the spectral signature of the electrodes, specifically the relative intensities of the sp2 band, sp3 line and sp2 to sp3 ratio, with measured electrical conductivity. In doing so, we demonstrate a practical pathway to in-line process monitoring and optimization for laser-written diamond electrodes.
II Results and Discussion
The laser-written graphitic pads and wires were chosen to emulate the two principal conductive elements found in devices. Figure 1a illustrates the experimental concept: a train of fs pulses (400 fs) is scanned through the surface of the diamond while a continuous-wave laser excited the same region and backscattered Raman emission is collected. Pads wider than 25 m and 200 m long reproduce the large‐area contact zones required for wire bonding or subsequent metallization, where sheet resistance and adhesion set the performance limit. By contrast, 1 m200 m wires replicate the sub-micron buried interconnects already deployed in high-density pixel detectors, biosensing micro-electrodes, and microwave striplines for NV-center control [9, 22, 23]. Studying both geometries therefore spans the full electrical landscape of laser-modified diamond; demonstrating a consistent correlation between Raman metrics and resistance across pads and wires is essential for future closed-loop fabrication. Accordingly, pads and wires of equal length were fabricated at varying laser scanning speeds to produce different degrees of graphitization, and thus a controlled range of electrical conductivities.
Figures 1b and c show optical micrographs along with photoluminescence (PL) and Raman maps (sp3 and sp2) for two laser-written graphitic wires fabricated at scan speeds of 20 and 200 m s-1. Of note, these maps were acquired post-fabrication, as complete datasets including the pristine diamond were required for the study. The wire written at 20 m s-1 exhibits a resistance of , whereas the 200 m s-1 wire shows , corresponding to an difference (see Figure 4b for complete set of conductivity measurements). Details of the graphitic pads are provided in Supplementary Note 1. Under optical illumination, the laser-written track appears dark relative to the surrounding single-crystal diamond. This contrast arises because conversion from transparent, wide-band-gap sp3 diamond to disordered sp2 carbon markedly increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [24, 25]. Therefore, the degree of darkening can be used as a first qualitative indication of graphitization. However, it conflates absorption, scattering, debris, and illumination artifacts, and it is not phase-specific. For this reason, PL and hyperspectral Raman mapping are regarded as more reliable approaches, since they provide phase-specific, quantitative, and spatially resolved indicators of graphitization.
In the PL maps (Figure 1b,c), the laser-written graphitized regions appear dark because the processing introduces non-radiative recombination pathways that quench the broadband fluorescence. While tracking broadband emission can provide valuable information, monitoring the depletion of the sp3 Raman peak yields higher spatial resolution and stronger phase-specific contrast between pristine diamond and graphitized areas. Indeed, local graphitization converts sp3–bonded diamond into sp2 carbon, depleting the sp3 signal and producing the pronounced contrast. Of note, the sp3 maps were constructed by integrating the hyperspectral signal across the first-order diamond peak (1325-1340cm-1) [24, 26], while the sp2 maps were obtained by integrating over 1575-1610 cm-1 [24]. In the latter, since graphitization leads to the formation of sp2-bonded carbon domains it results in a marked increase in the G-band intensity.
1 Spectral characterization of Laser-written Electrodes
As shown in Figure 1, one of the limitations of both brightfield microscopy and broadband PL imaging is that despite being good techniques for qualitative assessment of the fabrication they present limitations when used to provide quantitative information. As discussed in Supplementary Notes 3 and 4, PL could in principle provide quantitative evaluation of sp3 to sp2 conversion but, due to the broad integration, is very sensitive to other, unrelated phenomena, reducing quantitative fidelity. For this reason, we assessed the suitability of Raman spectroscopy for this task.
1.1 Single-band characterization
To determine how the degree of graphitization affects the sp3 to sp2 transformation, we performed hyperspectral Raman mapping on every graphitic wire and pad fabricated at distinct scan speeds (see Materials and Methods). Spectra belonging to the written features and to the surrounding pristine diamond were separated pixel-by-pixel and averaged. Figures 2a and b display representative spectra for (a) the unmodified single-crystal diamond and (b) a fully graphitized pad fabricated at a laser scan speed of 100 m s-1. The pristine crystal is dominated by the sp3 line at 1332 cm-1. In contrast, the graphitic structure written with low laser scan speed exhibits depleted sp3 intensity and is governed by the disorder-induced D band (1350 cm-1) together with the sp2 band (graphitic G band (1580 cm-1)), evidencing almost complete conversion to sp2 carbon.
Figure 2c compiles the speed-dependent averaged spectra of the laser-written graphitic wires. The spectral data normalized across all the samples to the sp3 peak (left panel) and the sp2 band (right panel), allowing the two regimes to be compared on equal footing across writing speeds. At the highest velocity tested (500 m s-1) the spectrum is sp3-dominated and the D and G bands are barely discernible, implying that only a small fraction of lattice sites have been converted to sp2. Progressively reducing the speed, and thereby increasing the deposited energy density per unit length, reverses this balance: the D and sp2 bands intensify while the sp3 line diminishes, signaling enhanced graphitization. The evolution is summarized quantitatively in Figure 2d, which plots the normalized integrated intensities and as well as their ratio versus laser scan speed. depletes with decreasing speed, whereas increases, except for a modest downturn below 50 m s-1, where excessive dose does not yields higher sp2 band intensity. The intensity ratio therefore spans almost a factor of three, from 0.9 at speeds 150 m s-1 to 0.3 at 500 m s-1, establishing laser scan speed as a single-parameter handle for tuning the local phase composition. A slight broadening of the sp3 line is also observed (see Supplementary Note 6), consistent with lattice strain introduced during partial sp2 formation.
1.2 Multi-band characterization
Laser-written tracks in diamond contain both unconverted sp3-bonded lattice and newly formed sp2-bonded carbon. Within a confocal voxel, these phases typically coexist as sub-micron domains, so a single Raman spectrum is an average over a heterogeneous local volume. In the previous section we showed that the degree of graphitization can be inferred from the intensity of the sp3 line. However, relying on a single peak can be fragile in practice because the apparent intensity can be influenced by various factors such as laser power fluctuations, defocusing, local topography and debris, baseline fluorescence, or strain-induced shifts/broadening.
To mitigate these effects, ratio metrics that combine several informative bands are usually preferred over single-peak readouts. In our case the most natural choice is to use the sp2/sp3 ratio, containing the information carried by the G/D complex and the 1332 cm-1 diamond line. Figure 3 illustrates this for a representative wire written at 50 m s-1: a 2D map of the ratio provides robust, self-normalized contrast across the field, enabling empirical thresholds that segment the scene into non-graphitized, partially graphitized, and fully graphitized classes. Class-averaged spectra make the assignments transparent: pristine regions show a narrow sp3 line accompanied by NV0 fluorescence, partially graphitized pixels exhibit concurrent sp3 attenuation and the emergence of D/G features, and fully graphitized regions are dominated by broad D/G bands with a strongly depleted 1332 cm-1 line.
While ratio maps remove much of the acquisition variability, they still require manual selection of specific bands. Hyperspectral unmixing offers a label-free alternative that exploits the entire spectrum and returns per-pixel composition in the form of abundances. As described in Materials and Methods, endmembers are extracted directly from the data using well established unmixing methods, and abundances are then estimated per pixel, which enables a direct fractional interpretation without the need for external references.
Applying this pipeline over a broad region of interest (1000–1800cm-1) yields endmembers that, while not identical to the actual reference spectra presented in Figure 2), preserve the expected relevant features: the diamond-associated endmember is characterised by a sharp 1332 cm-1 line and the NV0 contribution in the window, whereas the graphite-associated endmember exhibits a broad D/G complex with a suppressed diamond line. Importantly, this decomposition is entirely computational and label-free. The resulting abundance maps closely track the simpler sp3-window intensity maps, which could be anticipated given that the 1332 cm-1 line is the most intense and spatially specific feature, but provide a unified, quantitative representation across the whole spectrum.
To test sensitivity to the dominant diamond line, we repeated the analysis after excluding 1332 cm-1 (spectral ROI = 1345–1800 cm-1). In this case, the diamond-associated endmember is defined primarily by the NV0 contribution within the spectral ROI, while the complementary endmember remains the broad G/D complex. The corresponding abundance maps retain the expected spatial structure but exhibit lower signal-to-noise, as anticipated when the most intense, phase-specific line is removed. This cropped-ROI approach is nevertheless useful in scenarios where the diamond line saturates the detector, where strong local strain or temperature gradients shift or distort the 1332 cm-1 feature beyond a fixed window, or when one wishes to decouple the analysis from sp3-dominated contrast (e.g. to emphasize changes within the sp2 manifold).
Overall, the unmixing results reinforce the interpretation from ratio maps while removing the need to hand-pick bands: knowing only that two materials are present, the algorithm discovers representative endmembers and delivers per-pixel abundance maps that agree with the PL/Raman contrasts. These abundance maps offer a compact, quantitative metric for automated thresholding and closed-loop control (e.g. stop conditions based on sp3 depletion or sp2 abundance). Yet, in the particular case of monitoring graphitization in diamond, where the bands of interest are few and well known, hyperspectral unmixing does not provide a significant performance advantage over band-specific metrics.
2 Raman Monitoring of Electrical Performance
Ultimately, the optimal feedback signal for in-situ monitoring of fabrication should be directly tied to the functional performance of the resulting devices. For graphite electrodes on diamond, this performance is best represented by the conductivity of the laser-written features. To this end, the electrical conductivity of the pads and wires was measured using a custom setup (see Electrical conductivity measurement).
Figure 4 presents our findings and consolidates the link between laser scan speed, local bonding state, and device-level conductivity of the laser-written features. Figure 4a depicts the current–voltage traces for all wires and pads, which are linear over mV, indicating ohmic behavior and no measurable Schottky barriers; conduction is therefore governed by the volume fraction and connectivity of the laser-generated graphitic phase. For graphitic structure of width of 200 m written at ms-1, the pad resistance is , versus for the wire at the same speed. The lower pad resistance reflects its larger cross-section and additional over-writing during fabrication, which further graphitizes the diamond and increases the graphite fraction. Figure 4b generalises this comparison across scan speeds: the resistance decreases as the scan speed is reduced, with pad values consistently below wire values at identical speeds. All resistances are extracted from linear fits to quasi-static I–V sweeps within mV (details of the IV curves are shown in Supplementary Note 7).
In Figure 4c and d we plot, separately for wires and pads, the mean normalized integrated sp3 and sp2 intensities as well as their ratio versus laser scan speed. As the scan speed increases, the resistance rises concurrently with the normalized sp3 intensity, indicating reduced graphitization and weakened conductive pathways. For wires (Figure 4c), the normalized sp2 intensity increases significantly as the scan speed decreases from high to intermediate values, corresponding to a sharp drop in resistance. However, once the sp3 contribution is strongly diminished (deeply graphitized regime), sp2 intensity alone becomes a less reliable measure of further conductivity enhancement. Since the sp2 intensity rises as the sp3 intensity decreases with scan speed, their ratio mirrors these trends. Pads show the same qualitative behavior (Figure 4d) but remain systematically more conductive than wires at identical scan speeds, consistent with their larger effective cross-section and enhanced sp2 fraction. Collectively, these data establish Raman-derived sp2/sp3 metrics as a non-contact, quantitative proxy for electrical performance and a practical handle for in-situ process control during fs-laser fabrication.
3 Discussion
We have evaluated the performance of three characterization methods (optical micrography, PL mapping, and Raman microspectroscopy) for phase transformation from insulating sp3-bonded lattice to conductive sp2-bonded carbon, an ultrafast phase change that has been extensively studied in both surface and bulk geometries [26, 4, 27]. Among these characterization methods, only Raman provides a direct, quantitative, phase-specific measurement of the bonding state.
As shown in Figure 2, the sp3 and sp2 Raman features evolve with fabrication conditions and degree of graphitization. Reducing the scan speed, and thus increasing the energy deposited per unit length, leads to a steady decrease in sp3 intensity and a general increase in sp2 intensity, consistent with progressive phase conversion from diamond to graphitic carbon. At the slowest speeds, however, the sp2 intensity reaches a maximum and then declines despite continued sp3 depletion, a behavior also reported in other ultrafast graphitization studies[27, 28].
This counter-intuitive reduction in sp2 intensity can be interpreted within the ‘amorphization trajectory’ framework of Ferrari et al. [29, 30], wherein progressive disorder drives a transition from nanocrystalline graphite to amorphous sp2-rich carbon and, ultimately, to highly disordered or tetrahedral amorphous forms. In this regime, shrinking and fragmenting sp2 clusters suppress well-defined Raman modes, so the apparent sp2 signal diminishes even as the actual sp2 bonding fraction remains high. At intermediate doses, ordered sp2 clusters grow, strengthening the D and G bands. At extreme doses, however, high defect densities and reduced cluster sizes suppress well-defined vibrational modes, broadening and merging the D and G peaks into a low-contrast background. The resulting spectral degradation reduces the apparent sp2 Raman intensity, even when the actual sp2 bonding fraction remains high. Under such high-dose conditions, sp3 depletion provides a more reliable metric for tracking graphitization progression, as it decreases monotonically with increasing laser dose and is insensitive to spectral contrast loss[27, 25].
The stronger sp3 depletion observed for pads can be interpreted by stress-assisted graphitization during overwriting. Lattice mismatch between diamond and graphite generates tensile stress in the transformed surface layer and, crucially, in-plane compressive stress in the adjacent diamond [25]. Since the graphitization barrier decreases under lattice contraction, compression promotes earlier conversion as the lattice temperature rises. In pad writing, each new pass overlaps previously graphitized regions, introducing a local compressive field in neighboring diamond and enhancing the conversion rate. This local compressive strain field during overwriting explains the greater sp3 depletion and lower resistance of pads relative to single-pass wires at identical scan speeds, with the Raman shift direction matching established stress coefficients for diamond (blue shift in wavenumber) [31, 32, 33] (see Supplementary Note 6).
As presented in Figure 3, hyperspectral unmixing could also serve as a label-free, instrument-agnostic tool for generation of phase-fraction maps from 2D Raman data. The unmixing technique agrees with simpler ratio metrics while exploiting the full spectrum. By discovering two data-driven endmembers (ascribed to diamond and graphite phases) and enforcing non-negativity/sum-to-one constraints, the linear mixing model yields abundances that are directly interpretable as local composition and are robust to power drift, focus variations, and baseline fluorescence, issues that can affect single-band readouts. The approach is not without limits: severe disorder, stress-induced line shifts, or strong fluorescence can deviate from strict linearity and slightly bias abundance estimates.
Electrical measurements shown in Figure 4 reveal ohmic behavior for both pads and wires, with pads exhibiting systematically lower resistance due to their larger cross-section and cumulative overwriting during fabrication. Notably, the correlation between Raman metrics and resistance highlights a counter-intuitive result: sp3 depletion is a more robust indicator of electrical performance than the absolute sp2 signal. This distinction likely arises because sp3 suppression directly reflects the fraction of diamond lattice converted, irrespective of the ordering of the resulting carbon, whereas sp2 intensity is sensitive to disorder and may be suppressed even in highly conductive, amorphous-rich networks. Similar conclusions have been drawn in buried-wire studies where the attenuation of the diamond peak correlated linearly with conductive cross-section. Moreover, as shown in Figure 1, the superior SNR of the sp3 line relative to the sp2 band makes sp2-based metrics inherently more susceptible to broadband background from non-graphitic damage (see Supplementary Note 2). This is also true for full-spectrum based metrics, such as hyperspectral unmixing (see Supplementary Note 8).
From a metrology standpoint, this insight opens a streamlined path to in-situ control. In our experiments, one-second integrations of the diamond peak were sufficient to resolve sp3 depletion without sample cleaning. This can be further aided by the incidental removal of surface debris by the 532 nm probe laser (see Supplementary Note 1) which also reduces the masking effect of debris that might affect the fabrication.
III Conclusion and Outlook
Femtosecond laser writing in diamond produces conductive carbon features whose performance can be anticipated from phase-specific Raman spectral features. We have provided two-dimensional Raman mapping of laser-written surface structures under different fabrication conditions. Overall, our findings show that the sp2/sp3 intensity ratio provides a robust and reliable indicator of graphitization progression. Because it is derived from relative band intensities, this ratio is largely insensitive to variations in laser power and other experimental fluctuations. However, under high-dose irradiation conditions, monitoring sp3 depletion becomes a more consistent metric, as it decreases monotonically with increasing laser dose and remains unaffected by the loss of spectral contrast that accompanies extreme structural disorder [27, 25].
To make the readout robust, we complemented simple intensity ratios with blind hyperspectral unmixing performed without any pre-labelling. The algorithm discovers two endmembers (sp3-like and sp2-like) directly from the data and returns per-pixel abundances that agree with PL/Raman contrast while exploiting the full spectrum. This analysis also shows that comparing the background sp3 (pristine diamond) to the written sp2 provides the highest-SNR indicator of conversion and is therefore ideal for in-situ characterization. This means that a full spectrometer may not be necessary for feedback: the 1332 cm-1 line could be monitored using narrowband filtering and a CCD or APD, reducing system complexity and acquisition time. By establishing a depletion threshold corresponding to the target resistance, the fabrication process could be terminated or parameters adjusted in real time. Closed-loop optimization approaches, potentially incorporating machine learning for rapid spectral classification [34], could then maintain consistent electrical performance across large-area, multi-feature devices.
Beyond diamond and electrical performance, the same characteisation framework can be optimized and implemented in other materials by (a) acquiring (offline) representative sample datasets under the target fabrication conditions and apply the same label-free post-processing (ratio maps and blind unmixing) to discover robust, instrument-agnostic indicators and setpoints (e.g., depletion of a host phonon, bounds on D/G, endmember abundances); and (b) deploy those indicators as real-time monitors during fabrication using narrowband detection or lightweight spectral sampling. This strategy generalises as a local phase/composition sentinel across laser fabrication processes: (i) waveguide manufacturing in glasses and wide-bandgap crystals, where Raman-visible densification/strain fingerprints correlate with refractive-index change and loss [35, 36]; (ii) laser-induced graphene on polymers/resists [37], where D/G evolution guides sheet resistance and uniformity; (iii) embedded heaters/electrodes in SiC, sapphire, silica, and ceramics, where host phonons and carbon signatures define safe process windows [38, 39]; and (iv) phase-specific transformations in oxides and bio-relevant substrates, where a Raman marker provides a chemically selective stop condition [40].
In summary, hyperspectral Raman provides a phase‐specific, quantitative readout that links fabrication conditions to device performance during ultrafast laser writing in diamond. Across pads and wires, we find that depletion of the diamond sp3 line at 1332 cm-1 is a more reliable predictor of conductivity than absolute sp2 intensity, which can diminish at high disorder. A simple, self‐referenced sp3 metric, benchmarked to the local pristine background, therefore offers a robust in situ indicator, while a ratio-based metric, such as sp2-to-sp3 ratio offers better robustness to power, focus, and collection variations. Combined with label‐free unmixing to extract per‐pixel phase abundances in case of post-fabrication mapping, this framework enables calibrated thresholds for stop/adjust decisions and is readily implemented with narrowband detection for real‐time process control. Because the indicators are phase‐agnostic and instrument‐agnostic, the approach generalizes to other hosts and functionalities (e.g., waveguides, embedded heaters/electrodes, laser‐induced graphene), providing a practical route toward specification‐driven fs‐laser microfabrication.
IV Materials and Methods
1 Electrode fabrication
A single-crystal, optical-grade diamond grown by chemical-vapor deposition (Element Six; nominal N concentration ppm) served as the substrate for this study. Surface graphitic pads and wires were inscribed with a fs laser operating at 520 nm ( fs, kHz) with various laser scan speed. Wave-front aberrations were corrected by a liquid-crystal spatial light modulator, and the beam was focused through an objective lens with numerical aperture (ZEISS Plan-Neofluar, 0.5NA, 20x). Wire electrodes were written in a single pass under continuous stage translation (no overwriting). Successive pulses partially overlapped to form a continuous track; at fixed the pulse-to-pulse separation was set by the scan speed. Pad electrodes were produced by stacking parallel wires with a constant 0.5 m pitch, maintained for all scan speeds.
2 Electrical conductivity measurement
The resistance of laser-written pads and wires was determined from two-probe current–voltage (I–V) measurements using a low-noise transimpedance amplifier (DLPCA-200, FEMTO) together with a real-time voltage source and data-acquisition controller (ADwin-Gold II, Jäger). Electrical contact was made with two micromanipulated Pt probes. For pad devices, the probes contacted opposite ends of each laser-written pad. For wire devices, both termini were first connected to laser-written square pads (50 m 50 m, written at 100 m s-1), and the probes landed on these pads to standardize the contact area. The ADwin generated quasi-static voltage sweeps within the ohmic window (typically mV) and synchronously digitized the DLPCA-200 output.
3 Brightfield imaging
Optical micrographs were captured with an integrated transmission optical microscope. A red-LED array situated under the sample serves as illumination source and the images are captured on a CCD camera (EC650, Prosilica Inc.).
4 Confocal Imaging
Our fabrication setup has an integrated custom confocal microscope that can be simultaneously utilized for photoluminescence and Raman spectral collection. Using a 532 nm continuous-wave excitation laser (Cobolt Samba 150mW) and a fast steering mirror (FSM-300, Newport) relayed via a 4f into the back aperture of the system’s objective, 2D PL and RS maps can be collected. Notably, in this study, the 2D maps were collected post-fabrication utilising a higher numerical aperture objective (Olympus PlanApo 0.95NA, 60x).
Both photoluminescence and Raman spectra were collected simulataneously employing 30 mW of laser excitation. The samples were scanned with a 500nm pixel-to-pixel lateral displacement and a 1 second exposure time. The collected signal was split 99:1 between the Raman and the photoluminescence collection paths. For the latter, signal was collected via a single photon avalanche detector (SPCM-AQRH-14-FC, Excelitas Inc.). Given that the collection path of our confocal setup features a 550 nm long-pass dichroic, and that the SPAD quantum efficiency drops significantly after 850 nm, the PL collection window can be considered 550-850 nm. For the Raman collection path, scattered light was dispersed by a 1200 lines mm-1 blazed spectrograph (SpectraPro HRS-1200) and detected on a deep-cooled CCD camera (PIXIS 100, Princeton Instruments).
5 Raman spectra preprocessing
The hyperspectral Raman data were loaded as a three-dimensional array . To enable cross-dataset comparison (accounting for sample tilt, focus drift, and excitation-power fluctuations), we pre-normalized each map using pristine diamond regions identified from the sp3 window. Specifically, we integrated over cm-1 to obtain , ranked all pixels, and defined the brightest pixels (wire electrodes: ; pad electrodes: ) as the pristine reference. The normalization factor was the mean of these values, and the hyperspectral cube was normalized as , establishing .
From , we computed integrated intensities for sp3 (diamond, cm-1) and sp2 (graphitic, cm-1) at each pixel and identified darkest sp3 pixels ( for wire and for pad) as modified regions. For these pixels, we report the spatial coordinates, the normalized integrals and , their averages, and the ratio as a quantitative metric of graphitization. Fixed integration windows were used throughout (sp3 width cm-1; sp2 width cm-1); thus, the sp2/sp3 ratio is the ratio of integrated band intensities, not intensity per unit wavenumber. Full algorithmic details, parameter choices, and examples are provided in the SI (see Supplementary Note 3).
6 Spectral Unmixing
In this work we adopt the widely used linear mixing model (LMM), in which each spectrum is expressed as a linear combination of endmember spectra :
where denotes the abundance of the -th endmember, and .
Endmember identification.
To identify the endmember spectra from the data, we employed two well-established blind unmixing algorithms (a)Vertex Component Analysis (VCA) [41], a fast geometrical method that projects the dataset along random directions to iteratively identify extreme points of the spectral simplex, which serve as candidate endmembers; and (2) N-FINDR [42], an algorithm that searches for the set of endmembers maximising the volume of the simplex formed by candidate spectra, under the assumption that pure pixels are present in the dataset. Both algorithms require specifying the number of endmembers in advance. Here, we constrained the problem to two endmembers (sp2 and sp3), which reflects the known physics of the system and avoids overfitting.
Abundance estimation.
Once endmembers are identified, abundances can computed by solving the linear inverse problem under different constraints: (a) Fully Constrained Least Squares (FCLS) [43]: enforces both non-negativity (), and the sum-to-one constraint (), ensuring abundances can be interpreted as fractions; and (b) Non-Negative Least Squares (NNLS) [44]: relaxes the sum-to-one condition but maintains non-negativity, which can be useful when baseline variations are present. As shown in Figure 3, the abundance maps at 50 m/s, the pristine diamond region does not show an abundance of exactly 0 for the graphite endmember and 1 for the diamond endmember. This is expected since the endmembers are extracted across all wire maps, focus and intensity differences between datasets also shift the estimated simplex (see Supplementary Note 8). The result is a small residual assignment of sp2 abundance even in pristine regions. Importantly, the maps still provide clear contrast: pristine areas are dominated by the sp3 endmember, while graphitized tracks are dominated by sp2, confirming the physical validity of the decomposition.
Limits of linear mixing model.
Raman scattering is, to first order, linear in the number of scatterers, so spatial coexistence of sp3 and sp2 domains within the confocal volume leads to additive spectra, which the LMM captures. Moreover, our maps are recorded on polished surfaces with modest absorption at 532 nm outside the most heavily graphitized regions, reducing multiple scattering and re-absorption that could induce non-linear mixing. That said, three effects can challenge strict linearity: (i) strong broadband fluorescence in highly damaged zones, (ii) self-absorption and re-emission in thick, strongly absorbing tracks, and (iii) spectral distortions from stress/strain gradients that shift or broaden peaks in a way not representable by a fixed endmember. In such cases, non-linear or bilinear models (e.g., polynomial post–nonlinear mixing, kernel-based unmixing, or autoencoder-based unmixing [45]) may yield incremental gains, at the cost of additional parameters and reduced interpretability.
Implementation.
We conducted the hyperspectral unmixing analysis with Python using the RamanSPy package [46].
V Acknowledgments
This work was supported by the UK Engineering and Physical Sciences Research Council (EP/W025256/1).
VI Disclosures
The authors declare no conflicts of interest.
VII Data availability
The data that support the findings of this study are available from the corresponding authors, [X.C., A.F.G.], upon reasonable request.
Supplementary Information
Supplementary Note 1: Pad optical micrograph and Raman map
As desciribed in the manuscript, we employed laser-written graphitic pads and wires to emulate two common classes of conductive elements. Supplementary Figure 5a shows an optical micrograph of pad electrodes with a width of 50 m and length of 200 m, written at scan speeds of 100 and 200 m s-1, respectively. The 100 m s-1 pad exhibits a resistance of , whereas the 200 m s-1 pad shows , i.e., a difference of more than an order of magnitude. Under optical illumination, the laser-written tracks appear dark relative to the surrounding single-crystal diamond. This contrast arises because converting transparent, wide-band-gap sp3 diamond to disordered sp2 carbon increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [24, 25]. By contrast with the graphitic wires discussed in the main text, whose resistance difference is only weakly discernible in optical images, the two pads are optically indistinguishable. Supplementary Figures 5b and c present the corresponding photoluminescence (PL) and Raman (sp3 and sp2) maps. In both pads, the written regions appear bright in PL immediately after fabrication. This behavior differs from the wires, which appear dark in PL due to laser-induced non-radiative recombination pathways.
The bright pad emission consists of locally activated NV0 and NV- centres in this Type Ib diamond (high nitrogen content), and other defects centres that producing a broad band that photo-bleaches under green (532 nm) excitation. This highlights that monitoring the full-window PL signal is not an ideal metric during fabrication, as relevant phase information is obscured by unrelated emissions and does not directly reflect device performance. In contrast, hyperspectral Raman mapping provides more specific structural information. By integrating the sp3 diamond window (1325-1340cm-1) [26] and the sp2 G-band window (1575-1610cm-1) [24], the sp3 map shows the laser-written regions as decreased peak intensity relative to pristine diamond, with stronger depletion for the lower scan speed. The sp2 map exhibits, while exhibiting the expected behavior (i.e., increased peak intensity in the written region), has lower overall contrast. Both the sp3 and sp2 Raman maps exhibit horizontal line-like features, which arise from the writing process. Such features are more pronounced in the sp2 than in the sp3 maps. Details explanation will be discussed in Supplementary Note 2.
Supplementary Figure 6 shows an optical micrograph of a pad acquired after PL mapping. Scanning a 20 mW, 532 nm beam around the electrode removes the initial femtosecond-laser-induced graphite debris [25], revealing a clean structure (dashed green square) compared to the post-writing state (dashed red square). Thus, maintaining the 532 nm probe during Raman acquisition could serve the dual role of (a) enabling spectroscopy monitoring, and (b) clearing graphite debris. With this, debris would not mask the written features during in situ monitoring, and no additional cleaning step is required.
Supplementary Note 2: Origin of contrast in hyperspectral maps
As shown in Supplementary Figure 7, both the sp3 and sp2 Raman maps exhibit horizontal line-like features, which are more pronounced in the sp2 map than in the sp3 map. The origin of these contrasts differs for the two cases. Since the pad is fabricated by stacking parallel wires with a 0.5 m pitch, the bright features in the sp3 maps correspond to regions between adjacent wires where less damage occurs, resulting in reduced graphitization. In contrast, the bright features in the sp2 maps arise from a different mechanism, consistent with the behavior described in Figure 1c of the main text. Here, the increased brightness is not related to enhanced graphitization but instead originates from broadband background emission.
To confirm this, Supplementary Figure 7 presents a Raman map in the 590–600 nm spectral range (the same hyperspectral map of Figure 1c in the main text but different spectral window), excluding the sp3 and sp2 bands. Compared to the laser-written wire and the pure diamond background, the surrounding damaged shell appears bright, confirming that the observed contrast does not originate from either the sp3 or sp2 Raman signals. Spectra from three representative pixels—assigned to the graphite core, the damaged shell, and the pure diamond background—are extracted in Supplementary Figure 7a and shown in Supplementary Figure 7b–d. Each spectrum is self-normalized for comparison. The graphite region exhibits depleted sp3, a clear D band, and sp2 features, while the pure diamond spectrum is dominated by the sp3 line. In the damaged shell, no D or sp2 bands are observed apart from the sp3 peak. Instead, a broad emission is present, which contributes to the increased brightness observed in both the photoluminescence and hyperspectral maps. This emission is attributed to defect activation in the damaged shell surrounding the wire core [26], leading to an overall increase in count rates.
Supplementary Note 3: Pixel identification, integration, and normalization
The hyperspectral Raman data was loaded as a four-dimensional array , where and represent the spatial coordinates, denotes the focal depth, and represents the wavelength. For all analyses, we selected the -plane closest to µm, effectively reducing the data to a three-dimensional array . To compensate for experimental variations including focusing, and excitation laser power fluctuations between different datasets, we implemented a sample-specific normalization procedure based on pristine diamond regions (shown as green pixels in Fig. 8).
For the sp3 diamond peak analysis, we defined a wavelength window cm-1. The integrated intensity for each pixel was calculated as:
For every sample, we identified the brightest pixels (where for wire electrodes and for pad electrodes) by sorting all pixels according to their integrated intensity values. These brightest pixels correspond to pristine diamond regions that serve as our normalization baseline (see Fig. 8 where the green pixels indicate the normalization baseline). The normalization factor was computed as the mean intensity of these brightest pixels:
where represents the -th brightest pixel.
After establishing the normalization factor, the entire hyperspectral image was normalized:
Subsequently, we extracted and performed integrated intensity calculations for both the sp3 (diamond) and sp2 (graphitic carbon) spectral windows for those with dark sp3 intensity (red pixels in Fig. 8). For each pixel, we computed:
From the normalized sp3 integrated intensity map, we identified the darkest pixels (typically 60-2000 pixels depending on the sample types), which correspond to areas with reduced diamond features, indicating graphitization or structural modification.
The normalization procedure establishes a reference scale where a value of 1.0 represents the average integrated intensity of pristine diamond in the sp3 window (1325-1340 cm-1). Mathematically, . This normalization enables direct comparison between different datasets and quantitative assessment of the depletion of sp3 Raman signal. For the darkest pixels identified in our analysis, indicates reduced diamond sp3 character relative to pristine regions, while values can be directly compared to the sp3 baseline. The ratio provides a quantitative metric for the degree of graphitization, where indicates predominantly diamond sp3 character and suggests substantial sp2 carbon content.
For each identified dark pixel, we exported the normalized integrated intensities for both spectral windows along with their spatial coordinates. The average normalized intensity across all dark pixels was calculated as:
where represents the -th darkest pixel and . This averaging procedure provides a statistical measure of the overall modification in the analyzed region, with the normalized scale ensuring that values can be directly compared across different measurements and sample positions. The exported data enables quantitative comparison of the degree of graphitization across different electrode configurations and processing conditions. We note that the sp2 integration window (15 cm-1) is wider than the sp3 window (35 cm-1), reflecting the intrinsic spectral widths of these features in diamond. As all measurements use identical integration windows, this systematic difference does not affect relative comparisons between samples. The sp2/sp3 ratio should therefore be interpreted as the ratio of integrated intensities over their respective characteristic spectral features, rather than intensity per unit wavelength.
Supplementary Note 4: Quantitative PL analysis
To evaluate whether broadband photoluminescence (PL) intensity can serve as an in situ process indicator—analogous to the sp3, sp2 signals and their ratio, we extracted the PL intensity from the written pixels. The results are presented in Supplementary Figure 9. This figure shows a correlation between the resistance and the normalized PL intensity of the written wire electrodes, which suggests that broadband PL intensity can be leveraged to assess performance of the written electrodes.
However, given the broad spectral collection window of these PL maps (550-850 nm), they can suffer from significant fluctuations due to factors unrelated to the phenomenon of interest, in this case, graphitization. For example, in Type Ib diamond, however, intrinsic defects generate a non-negligible background in PL maps; consequently, the brightest and darkest PL pixels do not directly correspond to “pure diamond” or “graphitized” regions.
Supplementary Figure 9, in this particular case, since the sp3 component dominates the broadband PL for the written region, the normalized PL follows the sp3 trend: PL increases with resistance, consistent with a larger fraction of diamond-like bonding. Nevertheless, the broadband PL also includes emission from defect clusters, leading to higher uncertainty (i.e. measurement variance) than for the sp3 metric.
In summary, broadband PL can serve as a quantitative indicator during laser fabrication, but it is less specific and less accurate than the sp3 intensity. Accordingly, spectrally isolating the sp3 (and/or sp2) Raman windows—rather than relying on fully broadband PL—provides the most reliable and clean in situ process feedback.
Supplementary Note 5: sp3 distribution
The speed-dependent probability-density histograms of the normalised sp3 Raman intensity, computed from pixels identified as graphitised in the hyperspectral maps, are shown for pad and wire electrodes in Supplementary Figure 10 and Supplementary Figure 11, respectively. For each scan speed, 1500 pixels are analysed for pads and 60 for wires. Intensities are referenced to pristine diamond, so higher normalised sp3 corresponds to weaker graphitization.
In pad electrodes, increasing scan speed shifts the distributions toward higher values and simultaneously broadens them, consistent with reduced energy deposition per unit length and a less uniform sp3-to-sp2 transformation. By contrast, wire electrodes exhibit narrower distributions across all scan speeds because their geometry prevents overwriting. Nonetheless, they display the same overall trend as pads: increasing scan speed shifts the distributions to higher normalised sp3 values, reflecting a lower degree of graphitization in the laser-written structures.
Supplementary Note 6: sp3 Raman analysis
To complement the two-probe transport data, we quantified how the sp3 Raman response varies with laser scan speed. For each speed, 60 pixels within the graphitized region of a laser-written wire were averaged and the sp3 peak was fitted with a Lorentzian. The wire geometry was chosen over the pad because it provides a cleaner, more localized transformation without overlapping features. Averaged spectra for scan speeds from 10 to 1270 m s-1 were analyzed; the resulting peak center and full width at half maximum (FWHM) are plotted in Fig. 12 as blue circles and dark-orange squares, respectively. A pristine-diamond reference, obtained by averaging 100 pixels from an unexposed area, is shown at 3000m s-1 for visual comparison only.
Across the series, the sp3 mode lies slightly above 1332cm-1, consistent with tensile stress in the transformed graphite layer and the concomitant in-plane compressive stress imparted to the adjacent diamond[25], which produces a blue shift of the diamond Raman mode [31, 32]. The FWHM decreases systematically with increasing scan speed. Slower scans deposit more energy per unit length, driving a higher degree of graphitization; accordingly, the sp3 line departs from an ideal Lorentzian, broadening and sometimes becoming asymmetric. This behavior reflects increasing lattice disorder and strain distributions (inhomogeneous broadening), reduced phonon lifetimes (homogeneous broadening), and phonon-confinement effects; if the material becomes sufficiently conducting, Fano interference with an electronic continuum can further skew the lineshape [33, 47]. All averaged spectra and their Lorentzian fits are provided in Fig. 13.
Supplementary Note 7: Electrical performance for both pad and wire
Electrical measurements confirm that femtosecond-laser writing produces ohmic graphitic pathways whose conductance scales with scan speed and tracks the depletion of sp3 Raman intensity in the written region. For wire electrodes, the I-V characteristics are strictly linear from -0.5 to +0.5 V, and linear fits yield resistances that vary over several orders of magnitude as the scan speed increases from 10 to 1000 m s-1 (Fig. 14a-l). At the slowest scans (10-50 m s-1) the wires exhibit resistances in the tens of k, approaching the series resistance floor of the test board contact pads, whereas progressively faster scans—depositing less energy per unit length—produce markedly higher resistances reaching the M-G range.
Pad electrodes follow the same trend but with substantially lower absolute resistance due to their larger cross-section and multi-track overwriting, which yields more complete graphitization as indicated by the depleted sp3 Raman signal. I-V curves acquired from -0.1 to +0.1 V remain perfectly linear, with fitted resistances rising from a few hundred at 100 m s-1 to 50k at 500 m s-1 (Fig. 15a-e). The comparison between wires and pads highlights the expected geometric scaling and the writing process: wider, thicker pads support significantly higher conductance than narrow wires under otherwise identical conditions, and the zig-zag writing strategy introduces controlled overwriting that further enhances graphitization.
The electrical evolution with scan speed mirrors the Raman signatures of the laser-induced carbon. In both geometries, depletion of sp3 intensity and the concurrent strengthening of sp2 features at slower scan speeds coincide with reduced resistance, whereas the relative persistence of sp3 character at faster scans coincides with increased resistance. This direct Raman-electrical correlation substantiates that efficient graphitization—conversion from sp3 to sp2 bonding—governs charge transport in the written structures.
Supplementary Note 8: Hyperspectral unmixing
We performed hyperspectral unmixing with a linear mixing model using VCA for endmember extraction and FCLS for abundance estimation. As shown in Supplementary Figure 16, the unmixing results consistently show progressive loss of the sp3-like component with decreasing scan speed, while the sp2-like component increases and eventually saturates. This reinforces the main conclusion that sp3 depletion provides a reliable, monotonic indicator of graphitization. Hyperspectral unmixing is a powerful technique that leverages the full spectrum and does not require pre-defined labels. However, it is sensitive to the presence of noise or species not considered in the analysis. As shown in Supplementary Figure 16a, the presence of the broadband emission damage shell when fabricating at intermediate speeds is mislabeled as being a mixture of diamond and graphite phases.
Supplementary Figure 17 presents the endmembers extracted using different unmixing techniques and spectral regions of interest, exhibiting the variance that these endmembers can have. This behavior is expected: VCA and N-FINDR are geometric algorithms that approximate the data as lying within a convex simplex. They perform best when pixels lie close to the true extremes, but in practice the confocal voxel contains mixtures, variable SNR, NV0 fluorescence, and stress-induced peak shifts. The algorithms therefore identify different “extreme” representatives of the sp3-like and sp2-like responses, leading to small method-dependent variations that do not alter the physical interpretation. Indeed, the endmembers preserve the spectral features of their associated spectra, that is, the sp3 peak and NV0 band for the endmember associated to diamond, and the G band and, to lesser extent, the D band for the one associated to graphite.
Notably, for the ROI including the sp3 peak, the large dynamic range of this feature amplifies small fluctuations which can distort the convex geometry and bias endmember selection. Excluding this line (ROI 1345–1800 cm-1) reduces leverage and produces noisier but qualitatively similar maps.
It is important to distinguish geometric unmixing from PCA. VCA and N-FINDR seek spectra at the boundaries of the data cloud that can be interpreted as approximate endmembers, whereas PCA decomposes variance into orthogonal components that need not resemble real spectra and can take negative values. PCA is valuable for denoising and diagnostics, but its components cannot be interpreted as abundances.
(a) Spectral unmixing (VCA) on all wires.
(b) Spectral unmixing (VCA) on all pads.
(a) Endmembers when unmixing in the spectral ROI including sp3 (1000-1800 cm-1).
(b) Endmembers when unmixing in the spectral ROI not including sp3 (1345-1800 cm-1).

### v2

@@
 Abstract
-Raman microspectroscopy provides non-destructive, structure-specific contrast suited to in situ monitoring of femtosecond (fs) laser fabrication processes.
-In this work, we introduce a sample-agnostic Raman workflow that screens candidate spectral indicators using drift-resistant, self-normalized metrics and, where appropriate, label-free hyperspectral unmixing to extract endmembers and validate metric selection.
-Using diamond as an example, hyperspectral maps and electrical tests show that depletion of the 1332 cm-1 sp3 line is a monotonic, robust predictor of resistance—counterintuitively outperforming sp2-based measures often assumed optimal.
-This enables a practical implementation: a single narrowband channel at the selected marker provides real-time feedback without full-spectrum acquisition.
-Because the selection hinges on general band definitions and functional validation rather than diamond-specific physics, the approach translates Raman phase information directly into device specifications and is transferable to other Raman-active hosts for in-line quality control of laser-written structures.
-I Introduction
-Femtosecond laser writing enables direct, three-dimensional patterning of embedded features in transparent hosts via nonlinear absorption—multi-phonon ionization (MPI) and Zener breakdown—that initiates localized structural transformation [1, 2, 3].
-These processes permit buried and surface carbon wires and electrode architectures with arbitrary geometry for electrically conductive devices [4], including radiation detectors [5].
-Practical subsurface fabrication benefits from pre-compensating high-index–induced spherical aberration using adaptive optics such as a spatial light modulator (SLM) to restore a tight focal volume [6, 7].
-While the resulting microstructures have been examined by destructive or ex-situ methods—TEM [8, 9], SEM [2], and X-ray diffraction microscopy (XRD) [10]—their complexity and inability to operate in situ limit process monitoring.
-Raman microspectroscopy is a vibrational spectroscopic technique based on the inelastic scattering of monochromatic light.
-It has been extensively used for non-destructive identification, structural characterization, and monitoring of chemical and physical properties across materials science, chemistry, and biology[11].
-In carbon systems in particular, Raman can distinguish bonding configurations, crystallinity, and local strain, making it suitable for non-invasive, in situ, spatially resolved monitoring during fabrication [12].
-To obtain robust, comparable indicators from spectra acquired across fields of view and samples, previous studies have adopted a deconvolution and normalization strategy that favors self-referenced, phase-aware metrics over single-peak intensities [11].
-Absolute intensities are sensitive to power drift, focus/NA mismatch, baseline fluorescence, and strain; by contrast, combining phase-informative windows and leveraging label-free spectral deconvolution stabilizes the readout and enables cross-instrument comparison.
-In practice, we explicitly assign bands, normalize to pristine-material pixels, and compute complementary indicators of conversion (ratios and depletion measures defined below), converting hyperspectral maps into quantitative metrics suitable for in situ monitoring [12].
-Diamond is an instructive case for Raman-guided fs fabrication.
-As sp3-bonded carbon, it is an electrical insulator with exceptional hardness, high thermal conductivity, a wide bandgap, and chemical robustness, enabling applications in high-power electronics, photonics, and radiation detection.
-Point defects such as nitrogen vacancy centers further extend functionality to quantum information and nanoscale sensing [13, 14, 15, 16].
-The broader carbon design space—sp2-bonded forms (graphite, graphene, carbon nanotubes) with delocalized electrons for high electrical/thermal conductivity [17, 18], allotropic variants such as hexagonal diamond (lonsdaleite) [19], and mixed-phase nanocomposites such as ’diaphite’ [9, 20, 21]—underscores the value of a phase-resolved optical readout.
-Although diamond is transparent at 520 nm (2.38 eV) due to its 5.47 eV bandgap, which prevents linear photon absorption from the writing laser, the fabrication process instead relies on nonlinear absorption involving multi-phonon ionization (MPI) and Zener breakdown, initiating a confined sp3 to sp2 phase transition [1, 2, 3], forming local graphitized regions with device-relevant geometries.
-Here, we use Raman microspectroscopy to systematically assess fs-laser-written graphitic electrodes in diamond.
-Leveraging these vibrational modes, we perform direct mapping of surface graphitization through explicit band assignment and label-free spectral deconvolution, and we correlate the spectral signature of the electrodes, specifically the relative intensities of the sp2 band, sp3 line and sp2 to sp3 ratio, with measured electrical conductivity.
-In doing so, we demonstrate a practical pathway to in-line process monitoring and optimization for laser-written diamond electrodes.
-II Results and Discussion
-The laser-written graphitic pads and wires were chosen to emulate the two principal conductive elements found in devices.
-Figure 1a illustrates the experimental concept: a train of fs pulses (400 fs) is scanned through the surface of the diamond while a continuous-wave laser excited the same region and backscattered Raman emission is collected.
+Femtosecond laser fabrication enables the creation of a wide range of devices, but its scalability and yield can be limited by the lack of real-time, in-situ monitoring tools.
+In particular, there is a strong need for metrics that directly correlate with device performance.
+Raman microspectroscopy provides a non-destructive route for in-situ characterization.
+Here, we demonstrate its potential to assess the electrical performance of laser-written graphitic electrodes in diamond.
+By combining hyperspectral mapping with electrical testing, we show that depletion of the 1332 cm-1 sp3 Raman line serves as a monotonic and robust predictor of resistance, offering clear advantages over commonly used spectral features.
+We further introduce hyperspectral unmixing as a label-free approach to identify relevant spectral signatures in fabrication processes where Raman markers are less defined.
+Importantly, the methodology we present is not restricted to diamond but can be adapted to other host materials and functionalities, offering a practical path toward specification-driven fs-laser microfabrication.
+1 Department of Engineering Science, University of Oxford, Parks Road, Oxford OX1 3PJ, UK
+2 Department of Materials, University of Oxford, Parks Road, Oxford OX1 3PH, UK
+∗Corresponding authors: xingrui.cheng@eng.ox.ac.uk,
+alvaro.fernandezgaliana@eng.ox.ac.uk
+1 Introduction
+Femtosecond laser writing enables direct, three-dimensional patterning of embedded features in transparent hosts via nonlinear absorption—multi-phonon ionization and Zener breakdown—that initiates localized structural transformation [gattass2008femtosecond, musgraves2011laser].
+The technique has found particular application in processing of diamond [britel2025comparative, Jeschke1999microscopic, Ashikkalieva2016, Kononenko2016, sotillo2016], permitting fabrication of buried and surface carbon wires and electrode architectures with arbitrary geometry for electrically conductive devices [Sun2014], such as radiation detectors [komlenok2016diamond, Lagomarsino2013].
+Yet, one of the main current limitations for scaling manufacturing of these laser-written devices is the lack of feedback mechanisms that can be integrated for real-time, in-situ process monitoring and control.
+Indeed, most current structural analyses remain destructive or ex-situ (e.g.,TEM [Ashikkalieva2019, Salter2024], SEM [Ashikkalieva2016], and X-ray diffraction microscopy [Yin2024]).
+Raman microspectroscopy, based on inelastic light scattering, has been extensively used for non-destructive identification, structural characterization, and monitoring of chemical and physical properties across materials science, chemistry, and biology [fernandez2024fundamentals].
+In recent years, significant efforts have also focused on developing advanced computational approaches for spectral analysis, further enhancing and broadening its range of applications.
+Given its ability to differentiate the spectral signatures of sp3 versus sp2 bonding, crystallinity, and local strain environment, Raman spectroscopy is ideal for non-invasive, in-situ, spatially resolved characterization of carbon systems [khomich2022raman, Sotillo2018].
+In carbon systems, Raman signatures provide direct markers of bonding and structural order: graphitic sp2 domains give rise to the G band near , while disorder or finite crystallite size activates the D band at [Tuinstra1970, Ferrari2000].
+In contrast, single-crystal diamond exhibits a sharp first-order mode at [Knight1989, Ferrari2004].
+Throughout this work, we denote these features as the sp3 line (), sp2 band (), and D band (), providing a consistent basis for assessing carbon phase composition, crystallinity, and defect populations in laser-written structures.
+Here, we use Raman microspectroscopy to systematically assess fs-laser-written graphitic electrodes in diamond, and compare its performance with brightfield and photoluminescence imaging.
+Leveraging phase-specific vibrational modes of carbon, we perform direct mapping of surface graphitization through explicit band assignment and label-free spectral deconvolution, and correlate the spectral signatures to the device performance (i.e., electrical conductivity).
+Beyond offering detailed insights into graphitization monitoring in diamond, the methodology outlined here is broadly transferable: it can be applied to evaluate Raman spectroscopy as a performance probe in diverse device platforms, since none of its principles are inherently diamond-specific.
+2 Results and Discussion
+To model the principal conductive elements in diamond devices, we use a femtosecond laser to fabricate graphitic pads and wires on the surface of a CVD diamond sample, as shown in Figure 1a.
 Pads wider than 25 m and 200 m long reproduce the large‐area contact zones required for wire bonding or subsequent metallization, where sheet resistance and adhesion set the performance limit.
-By contrast, 1 m200 m wires replicate the sub-micron buried interconnects already deployed in high-density pixel detectors, biosensing micro-electrodes, and microwave striplines for NV-center control [9, 22, 23].
-Studying both geometries therefore spans the full electrical landscape of laser-modified diamond; demonstrating a consistent correlation between Raman metrics and resistance across pads and wires is essential for future closed-loop fabrication.
-Accordingly, pads and wires of equal length were fabricated at varying laser scanning speeds to produce different degrees of graphitization, and thus a controlled range of electrical conductivities.
-Figures 1b and c show optical micrographs along with photoluminescence (PL) and Raman maps (sp3 and sp2) for two laser-written graphitic wires fabricated at scan speeds of 20 and 200 m s-1.
-Of note, these maps were acquired post-fabrication, as complete datasets including the pristine diamond were required for the study.
+By contrast, 1 m 200 m wires replicate the sub-micron buried interconnects already deployed in high-density pixel detectors, biosensing micro-electrodes, and microwave striplines for NV-center control [Salter2024, Bloomer2020, FORNERIS2017].
+Studying both geometries spans the full electrical landscape of laser-modified diamond.
+Moreover, pads and wires of equal length were fabricated at varying laser scanning speeds to produce different degrees of graphitization, and thus a controlled range of electrical conductivities.
+For each of these structures, brightfield transmission microscopy images were captured, along with photoluminescence (PL) and backscattered Raman, collected using our integrated confocal microscope (see Materials and Methods).
+Figures 1b and c show these optical micrographs along with PL and Raman maps (sp3 and sp2) for two laser-written graphitic wires fabricated at scan speeds of 20 and 200 m s-1.
 The wire written at 20 m s-1 exhibits a resistance of , whereas the 200 m s-1 wire shows , corresponding to an difference (see Figure 4b for complete set of conductivity measurements).
-Details of the graphitic pads are provided in Supplementary Note 1.
+Details of the graphitic pads are provided in Supporting Note 1.
 Under optical illumination, the laser-written track appears dark relative to the surrounding single-crystal diamond.
-This contrast arises because conversion from transparent, wide-band-gap sp3 diamond to disordered sp2 carbon markedly increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [24, 25].
+This contrast arises because conversion from transparent, wide-band-gap sp3 diamond to disordered sp2 carbon markedly increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [prelas1997handbook, Wang2000].
 Therefore, the degree of darkening can be used as a first qualitative indication of graphitization.
@@
 In the PL maps (Figure 1b,c), the laser-written graphitized regions appear dark because the processing introduces non-radiative recombination pathways that quench the broadband fluorescence.
-While tracking broadband emission can provide valuable information, monitoring the depletion of the sp3 Raman peak yields higher spatial resolution and stronger phase-specific contrast between pristine diamond and graphitized areas.
-Indeed, local graphitization converts sp3–bonded diamond into sp2 carbon, depleting the sp3 signal and producing the pronounced contrast.
-Of note, the sp3 maps were constructed by integrating the hyperspectral signal across the first-order diamond peak (1325-1340cm-1) [24, 26], while the sp2 maps were obtained by integrating over 1575-1610 cm-1 [24].
-In the latter, since graphitization leads to the formation of sp2-bonded carbon domains it results in a marked increase in the G-band intensity.
-1 Spectral characterization of Laser-written Electrodes
-As shown in Figure 1, one of the limitations of both brightfield microscopy and broadband PL imaging is that despite being good techniques for qualitative assessment of the fabrication they present limitations when used to provide quantitative information.
-As discussed in Supplementary Notes 3 and 4, PL could in principle provide quantitative evaluation of sp3 to sp2 conversion but, due to the broad integration, is very sensitive to other, unrelated phenomena, reducing quantitative fidelity.
-For this reason, we assessed the suitability of Raman spectroscopy for this task.
-1.1 Single-band characterization
-To determine how the degree of graphitization affects the sp3 to sp2 transformation, we performed hyperspectral Raman mapping on every graphitic wire and pad fabricated at distinct scan speeds (see Materials and Methods).
-Spectra belonging to the written features and to the surrounding pristine diamond were separated pixel-by-pixel and averaged.
+Yet, it is not phase-specific and can be easily affected by broadband emission, as shown in Figure 1c and detailed in Supporting Note 2.
+On the other hand, Raman mapping provides phase-specific contrast.
+Local graphitization converts sp3 diamond to sp2 carbon, depleting the sp3 line and enhancing the G band; accordingly, we construct sp3 maps by integrating 1325–1340 cm-1 and sp2 maps by integrating 1575–1610 cm-1 [prelas1997handbook, Kononenko2008].
+These phase-specific features (sp3 line, sp2 band, D band) enable quantitative, cross-structure comparison of phase content, crystallinity, and defect populations in laser-written carbon [Sotillo2018, Kononenko2008, Ashikkalieva2019].
+Note that, as shown in (Figure 1c), Raman signal can also be contaminated by non-specific fluorescent background, an effect that can oftentimes be minimized with baseline correction and other techniques [wei2015review].
+2.1 Spectral characterization of Laser-written Electrodes
+As shown in Figure 1, one of the limitations of both brightfield microscopy and broadband PL is that despite being good techniques for qualitative assessment of the fabrication they present limitations when used to provide phase-specific quantitative information.
+In the case of PL, it can in principle be used for quantitative sp3-to-sp2 conversion evaluation, but its sensitivity to broadband emission not stemming from the process of interest reduces its fidelity.
+This is further discussed in Supporting Notes 3 and 4.
+Therefore, we assessed the suitability of Raman spectroscopy for this task.
+In graphitic carbons, the in-plane first-order phonon (“G band”) appears near , while disorder or finite crystallite size activates the defect-induced “D band” around [Tuinstra1970, Ferrari2000].
+In single-crystal diamond, the zone-center optical phonon of symmetry gives a sharp first-order line at [Knight1989, Ferrari2004].
+These emphasize the phase origin of each peak and facilitates quantitative comparison across laser-written structures, enabling identification and quantification of carbon phases, crystallinity, and defect populations [Sotillo2018, Kononenko2008, Ashikkalieva2019].
+2.1.1 Single-band characterization
+Using the integrated confocal microscope, we collected hyperspectral Raman maps on every graphitic wire and pad fabricated at distinct scan speeds (see Materials and Methods).
 Figures 2a and b display representative spectra for (a) the unmodified single-crystal diamond and (b) a fully graphitized pad fabricated at a laser scan speed of 100 m s-1.
@@
 The evolution is summarized quantitatively in Figure 2d, which plots the normalized integrated intensities and as well as their ratio versus laser scan speed.
-depletes with decreasing speed, whereas increases, except for a modest downturn below 50 m s-1, where excessive dose does not yields higher sp2 band intensity.
-The intensity ratio therefore spans almost a factor of three, from 0.9 at speeds 150 m s-1 to 0.3 at 500 m s-1, establishing laser scan speed as a single-parameter handle for tuning the local phase composition.
-A slight broadening of the sp3 line is also observed (see Supplementary Note 6), consistent with lattice strain introduced during partial sp2 formation.
-1.2 Multi-band characterization
-Laser-written tracks in diamond contain both unconverted sp3-bonded lattice and newly formed sp2-bonded carbon.
+depletes with decreasing speed, whereas increases, except for a modest downturn below 50 m s-1, where excessive dose does not yield higher sp2 band intensity.
+2.1.2 Multi-band characterization
+Laser-written electrodes in diamond contain both unconverted sp3-bonded lattice and newly formed sp2-bonded carbon.
 Within a confocal voxel, these phases typically coexist as sub-micron domains, so a single Raman spectrum is an average over a heterogeneous local volume.
-In the previous section we showed that the degree of graphitization can be inferred from the intensity of the sp3 line.
+In the previous section we showed that the degree of graphitization can be inferred from the intensity of certain bands (e.g., sp3 line).
 However, relying on a single peak can be fragile in practice because the apparent intensity can be influenced by various factors such as laser power fluctuations, defocusing, local topography and debris, baseline fluorescence, or strain-induced shifts/broadening.
@@
 In our case the most natural choice is to use the sp2/sp3 ratio, containing the information carried by the G/D complex and the 1332 cm-1 diamond line.
-Figure 3 illustrates this for a representative wire written at 50 m s-1: a 2D map of the ratio provides robust, self-normalized contrast across the field, enabling empirical thresholds that segment the scene into non-graphitized, partially graphitized, and fully graphitized classes.
-Class-averaged spectra make the assignments transparent: pristine regions show a narrow sp3 line accompanied by NV0 fluorescence, partially graphitized pixels exhibit concurrent sp3 attenuation and the emergence of D/G features, and fully graphitized regions are dominated by broad D/G bands with a strongly depleted 1332 cm-1 line.
+Figure 3 illustrates this for a representative wire written at 50 m s-1.
+2.1.3 Hyperspectral unmixing
 While ratio maps remove much of the acquisition variability, they still require manual selection of specific bands.
-Hyperspectral unmixing offers a label-free alternative that exploits the entire spectrum and returns per-pixel composition in the form of abundances.
-As described in Materials and Methods, endmembers are extracted directly from the data using well established unmixing methods, and abundances are then estimated per pixel, which enables a direct fractional interpretation without the need for external references.
-Applying this pipeline over a broad region of interest (1000–1800cm-1) yields endmembers that, while not identical to the actual reference spectra presented in Figure 2), preserve the expected relevant features: the diamond-associated endmember is characterised by a sharp 1332 cm-1 line and the NV0 contribution in the window, whereas the graphite-associated endmember exhibits a broad D/G complex with a suppressed diamond line.
-Importantly, this decomposition is entirely computational and label-free.
-The resulting abundance maps closely track the simpler sp3-window intensity maps, which could be anticipated given that the 1332 cm-1 line is the most intense and spatially specific feature, but provide a unified, quantitative representation across the whole spectrum.
+Therefore, we also introduce the use of hyperspectral unmixing as a label-free alternative for fabrication processes where the bands of each phase might not be known.
+As described in Materials and Methods, we model each spectrum as a linear mixture of endmembers and use well-established linear unmixing methods, such as vector component analysis (VCA) [nascimento2005vertex] to estimate these endmembers directly from the data.
+Then, abundances are estimated per pixel solving the linear inverse problem.
+In particular, for the results shown in Figure 3, we impose endmembers and enforce both non-negativity, and the sum-to-one of the abundances at every pixel using fully constrained least squares (FCLS) [heinz2001fully], which enables a direct fractional interpretation without the need for external references.
+As shown in Figure 3e, when applied over a broad region of interest (1000–1800cm-1) unmixing yields endmembers that, while not identical to the actual reference spectra presented in Figure 3), preserve the expected relevant features: the diamond-associated endmember is characterised by a sharp 1332 cm-1 line, whereas the graphite-associated endmember exhibits a broad D/G complex with a suppressed diamond line.
+Additionally, the diamond-associated endmember also features a broader band centered around cm-1, which is consistent with the zero-phonon line of neutral single nitrogen-vacancy defects (NV0) in diamond [Manson2013].
+The resulting abundance maps also closely track the simpler sp3-window intensity maps (Figure 3b).
 To test sensitivity to the dominant diamond line, we repeated the analysis after excluding 1332 cm-1 (spectral ROI = 1345–1800 cm-1).
 In this case, the diamond-associated endmember is defined primarily by the NV0 contribution within the spectral ROI, while the complementary endmember remains the broad G/D complex.
-The corresponding abundance maps retain the expected spatial structure but exhibit lower signal-to-noise, as anticipated when the most intense, phase-specific line is removed.
-This cropped-ROI approach is nevertheless useful in scenarios where the diamond line saturates the detector, where strong local strain or temperature gradients shift or distort the 1332 cm-1 feature beyond a fixed window, or when one wishes to decouple the analysis from sp3-dominated contrast (e.g.
-to emphasize changes within the sp2 manifold).
+The corresponding abundance maps retain the expected spatial structure but exhibit lower signal-to-noise (SNR), as anticipated when the most intense, phase-specific line is removed.
 Overall, the unmixing results reinforce the interpretation from ratio maps while removing the need to hand-pick bands: knowing only that two materials are present, the algorithm discovers representative endmembers and delivers per-pixel abundance maps that agree with the PL/Raman contrasts.
-These abundance maps offer a compact, quantitative metric for automated thresholding and closed-loop control (e.g.
-stop conditions based on sp3 depletion or sp2 abundance).
-Yet, in the particular case of monitoring graphitization in diamond, where the bands of interest are few and well known, hyperspectral unmixing does not provide a significant performance advantage over band-specific metrics.
-2 Raman Monitoring of Electrical Performance
+2.2 Raman Monitoring of Electrical Performance
 Ultimately, the optimal feedback signal for in-situ monitoring of fabrication should be directly tied to the functional performance of the resulting devices.
 For graphite electrodes on diamond, this performance is best represented by the conductivity of the laser-written features.
-To this end, the electrical conductivity of the pads and wires was measured using a custom setup (see Electrical conductivity measurement).
+To this end, the electrical conductivity of the pads and wires was measured using a custom setup (see Materials and Methods).
 Figure 4 presents our findings and consolidates the link between laser scan speed, local bonding state, and device-level conductivity of the laser-written features.
@@
 The lower pad resistance reflects its larger cross-section and additional over-writing during fabrication, which further graphitizes the diamond and increases the graphite fraction.
-Figure 4b generalises this comparison across scan speeds: the resistance decreases as the scan speed is reduced, with pad values consistently below wire values at identical speeds.
-All resistances are extracted from linear fits to quasi-static I–V sweeps within mV (details of the IV curves are shown in Supplementary Note 7).
+All resistances are extracted from linear fits to quasi-static I–V sweeps within mV (details of the IV curves are shown in Supporting Note 10 and 11).
 In Figure 4c and d we plot, separately for wires and pads, the mean normalized integrated sp3 and sp2 intensities as well as their ratio versus laser scan speed.
 As the scan speed increases, the resistance rises concurrently with the normalized sp3 intensity, indicating reduced graphitization and weakened conductive pathways.
-For wires (Figure 4c), the normalized sp2 intensity increases significantly as the scan speed decreases from high to intermediate values, corresponding to a sharp drop in resistance.
+For wires (Fig.
+4c), decreasing the scan speed from high to intermediate increases the normalized sp2 intensity and lowers the resistance.
 However, once the sp3 contribution is strongly diminished (deeply graphitized regime), sp2 intensity alone becomes a less reliable measure of further conductivity enhancement.
 Since the sp2 intensity rises as the sp3 intensity decreases with scan speed, their ratio mirrors these trends.
-Pads show the same qualitative behavior (Figure 4d) but remain systematically more conductive than wires at identical scan speeds, consistent with their larger effective cross-section and enhanced sp2 fraction.
-Collectively, these data establish Raman-derived sp2/sp3 metrics as a non-contact, quantitative proxy for electrical performance and a practical handle for in-situ process control during fs-laser fabrication.
-3 Discussion
-We have evaluated the performance of three characterization methods (optical micrography, PL mapping, and Raman microspectroscopy) for phase transformation from insulating sp3-bonded lattice to conductive sp2-bonded carbon, an ultrafast phase change that has been extensively studied in both surface and bulk geometries [26, 4, 27].
-Among these characterization methods, only Raman provides a direct, quantitative, phase-specific measurement of the bonding state.
-As shown in Figure 2, the sp3 and sp2 Raman features evolve with fabrication conditions and degree of graphitization.
-Reducing the scan speed, and thus increasing the energy deposited per unit length, leads to a steady decrease in sp3 intensity and a general increase in sp2 intensity, consistent with progressive phase conversion from diamond to graphitic carbon.
-At the slowest speeds, however, the sp2 intensity reaches a maximum and then declines despite continued sp3 depletion, a behavior also reported in other ultrafast graphitization studies[27, 28].
+Pads show the same qualitative behavior (Figure 4d) but remain systematically more conductive than wires at identical scan speeds.
+2.3 Discussion
+Figure 2 summarizes how the Raman features evolve with writing dose, controlled by scan speed.
+Reducing the speed, which increases the energy deposited per unit length, produces a monotonic depletion of the 1332 cm-1 sp3 line and a general rise of the 1580 cm-1 sp2 band, consistent with progressive diamond-to-graphite conversion.
+At the slowest speeds, however, the sp2 intensity reaches a maximum and then declines despite continued sp3 depletion, a behavior also reported in other studies on diamond graphitization [Ali2022, Khomich2023].
 This counter-intuitive reduction in sp2 intensity can be interpreted within the ‘amorphization trajectory’ framework of Ferrari et al.
-[29, 30], wherein progressive disorder drives a transition from nanocrystalline graphite to amorphous sp2-rich carbon and, ultimately, to highly disordered or tetrahedral amorphous forms.
+[Ferrari2000, Ferrari2004], wherein progressive disorder drives a transition from nanocrystalline graphite to amorphous sp2-rich carbon and, ultimately, to highly disordered or tetrahedral amorphous forms.
 In this regime, shrinking and fragmenting sp2 clusters suppress well-defined Raman modes, so the apparent sp2 signal diminishes even as the actual sp2 bonding fraction remains high.
@@
 The resulting spectral degradation reduces the apparent sp2 Raman intensity, even when the actual sp2 bonding fraction remains high.
-Under such high-dose conditions, sp3 depletion provides a more reliable metric for tracking graphitization progression, as it decreases monotonically with increasing laser dose and is insensitive to spectral contrast loss[27, 25].
-The stronger sp3 depletion observed for pads can be interpreted by stress-assisted graphitization during overwriting.
-Lattice mismatch between diamond and graphite generates tensile stress in the transformed surface layer and, crucially, in-plane compressive stress in the adjacent diamond [25].
-Since the graphitization barrier decreases under lattice contraction, compression promotes earlier conversion as the lattice temperature rises.
-In pad writing, each new pass overlaps previously graphitized regions, introducing a local compressive field in neighboring diamond and enhancing the conversion rate.
-This local compressive strain field during overwriting explains the greater sp3 depletion and lower resistance of pads relative to single-pass wires at identical scan speeds, with the Raman shift direction matching established stress coefficients for diamond (blue shift in wavenumber) [31, 32, 33] (see Supplementary Note 6).
-As presented in Figure 3, hyperspectral unmixing could also serve as a label-free, instrument-agnostic tool for generation of phase-fraction maps from 2D Raman data.
-The unmixing technique agrees with simpler ratio metrics while exploiting the full spectrum.
-By discovering two data-driven endmembers (ascribed to diamond and graphite phases) and enforcing non-negativity/sum-to-one constraints, the linear mixing model yields abundances that are directly interpretable as local composition and are robust to power drift, focus variations, and baseline fluorescence, issues that can affect single-band readouts.
-The approach is not without limits: severe disorder, stress-induced line shifts, or strong fluorescence can deviate from strict linearity and slightly bias abundance estimates.
-Electrical measurements shown in Figure 4 reveal ohmic behavior for both pads and wires, with pads exhibiting systematically lower resistance due to their larger cross-section and cumulative overwriting during fabrication.
-Notably, the correlation between Raman metrics and resistance highlights a counter-intuitive result: sp3 depletion is a more robust indicator of electrical performance than the absolute sp2 signal.
-This distinction likely arises because sp3 suppression directly reflects the fraction of diamond lattice converted, irrespective of the ordering of the resulting carbon, whereas sp2 intensity is sensitive to disorder and may be suppressed even in highly conductive, amorphous-rich networks.
-Similar conclusions have been drawn in buried-wire studies where the attenuation of the diamond peak correlated linearly with conductive cross-section.
-Moreover, as shown in Figure 1, the superior SNR of the sp3 line relative to the sp2 band makes sp2-based metrics inherently more susceptible to broadband background from non-graphitic damage (see Supplementary Note 2).
-This is also true for full-spectrum based metrics, such as hyperspectral unmixing (see Supplementary Note 8).
-From a metrology standpoint, this insight opens a streamlined path to in-situ control.
-In our experiments, one-second integrations of the diamond peak were sufficient to resolve sp3 depletion without sample cleaning.
-This can be further aided by the incidental removal of surface debris by the 532 nm probe laser (see Supplementary Note 1) which also reduces the masking effect of debris that might affect the fabrication.
-III Conclusion and Outlook
-Femtosecond laser writing in diamond produces conductive carbon features whose performance can be anticipated from phase-specific Raman spectral features.
-We have provided two-dimensional Raman mapping of laser-written surface structures under different fabrication conditions.
-Overall, our findings show that the sp2/sp3 intensity ratio provides a robust and reliable indicator of graphitization progression.
-Because it is derived from relative band intensities, this ratio is largely insensitive to variations in laser power and other experimental fluctuations.
-However, under high-dose irradiation conditions, monitoring sp3 depletion becomes a more consistent metric, as it decreases monotonically with increasing laser dose and remains unaffected by the loss of spectral contrast that accompanies extreme structural disorder [27, 25].
-To make the readout robust, we complemented simple intensity ratios with blind hyperspectral unmixing performed without any pre-labelling.
-The algorithm discovers two endmembers (sp3-like and sp2-like) directly from the data and returns per-pixel abundances that agree with PL/Raman contrast while exploiting the full spectrum.
-This analysis also shows that comparing the background sp3 (pristine diamond) to the written sp2 provides the highest-SNR indicator of conversion and is therefore ideal for in-situ characterization.
-This means that a full spectrometer may not be necessary for feedback: the 1332 cm-1 line could be monitored using narrowband filtering and a CCD or APD, reducing system complexity and acquisition time.
-By establishing a depletion threshold corresponding to the target resistance, the fabrication process could be terminated or parameters adjusted in real time.
-Closed-loop optimization approaches, potentially incorporating machine learning for rapid spectral classification [34], could then maintain consistent electrical performance across large-area, multi-feature devices.
-Beyond diamond and electrical performance, the same characteisation framework can be optimized and implemented in other materials by (a) acquiring (offline) representative sample datasets under the target fabrication conditions and apply the same label-free post-processing (ratio maps and blind unmixing) to discover robust, instrument-agnostic indicators and setpoints (e.g., depletion of a host phonon, bounds on D/G, endmember abundances); and (b) deploy those indicators as real-time monitors during fabrication using narrowband detection or lightweight spectral sampling.
-This strategy generalises as a local phase/composition sentinel across laser fabrication processes: (i) waveguide manufacturing in glasses and wide-bandgap crystals, where Raman-visible densification/strain fingerprints correlate with refractive-index change and loss [35, 36]; (ii) laser-induced graphene on polymers/resists [37], where D/G evolution guides sheet resistance and uniformity; (iii) embedded heaters/electrodes in SiC, sapphire, silica, and ceramics, where host phonons and carbon signatures define safe process windows [38, 39]; and (iv) phase-specific transformations in oxides and bio-relevant substrates, where a Raman marker provides a chemically selective stop condition [40].
-In summary, hyperspectral Raman provides a phase‐specific, quantitative readout that links fabrication conditions to device performance during ultrafast laser writing in diamond.
-Across pads and wires, we find that depletion of the diamond sp3 line at 1332 cm-1 is a more reliable predictor of conductivity than absolute sp2 intensity, which can diminish at high disorder.
-A simple, self‐referenced sp3 metric, benchmarked to the local pristine background, therefore offers a robust in situ indicator, while a ratio-based metric, such as sp2-to-sp3 ratio offers better robustness to power, focus, and collection variations.
-Combined with label‐free unmixing to extract per‐pixel phase abundances in case of post-fabrication mapping, this framework enables calibrated thresholds for stop/adjust decisions and is readily implemented with narrowband detection for real‐time process control.
-Because the indicators are phase‐agnostic and instrument‐agnostic, the approach generalizes to other hosts and functionalities (e.g., waveguides, embedded heaters/electrodes, laser‐induced graphene), providing a practical route toward specification‐driven fs‐laser microfabrication.
-IV Materials and Methods
-1 Electrode fabrication
+Therefore, under such high-dose conditions, sp3 depletion provides a more reliable metric for tracking graphitization progression, as it decreases monotonically with increasing laser dose, is insensitive to spectral contrast loss, and, thanks to its higher SNR, is less sensitive to non-specific broadband background [Ali2022, Wang2000].
+Pad writing shows stronger sp3 depletion than single-pass wires, which we attribute to stress-assisted graphitization during overwriting.
+Lattice mismatch generates tensile stress in the transformed layer and in-plane compressive stress in adjacent diamond [Wang2000].
+Compression lowers the graphitization barrier, so overlapping passes enhance conversion.
+The observed wavenumber shift of the diamond line is consistent with established stress coefficients [Grimsditch1975, Zaitsev2001, Orwa2000] (see Supporting Note 6) and explains the lower resistance measured for pads at identical scan speeds.
+In Figure 3 we introduce the use of more advanced Raman-based structure-characterization metrics.
+Particularly we first present multi-band metrics, mainly sp3-to-sp2 ratio.
+Using this ratio has the advantage of being agnostic to some of the effects that can generate non-structure-specific intensity fluctuations in the sp2 and sp3 peaks, such as excitation power.
+Moreover, we also introduce hyperspectral unmixing as a label-free complementary spectral analysis (Fig.
+3).
+A linear mixing model with non-negativity and sum-to-one constraints recovers two data-driven endmembers that map onto diamond-like and graphite-like responses, producing phase-fraction maps that agree with band-based indicators while leveraging the full spectrum.
+The approach is advantageous as it does not rely on pre-established bands of interest and is also robust to power drift and can potentially help reduce the effect of baseline fluorescence, though the linearity assumption presents some limitations (see Materials and Methods).
+Yet, in the case of monitoring graphitization in diamond, where the bands of interest are well-known, hyperspectral unmixing does not provide a significant advantage.
+Finally, in Figure 4 we show how Raman spectroscopy can be used to evaluate the performance of graphite electrodes.
+Across devices, depletion of the 1332 cm-1 sp3 line correlates with resistance more robustly than absolute sp2 intensity.
+This is likely due to a combination of the sensitivity of sp2 to disorder in high-dose conditions and the effect of the broadband background from nongraphitic damage, which is more severe in sp2 in the low-dose regime due to the difference in SNR with the sp3 line (Fig.
+1, Supporting Note 2).
+As shown in Figures 4c,d, the sp2-to-sp3 ratio is a useful indicator of graphitization because relative normalization reduces sensitivity to power fluctuations.
+Yet, under high-dose irradiation, sp3 exhibits less variance than sp2-to-sp3 ratio, likely due to the effect of extreme disorder on the sp2 band [Ali2022, Wang2000].
+3 Conclusion and Outlook
+Raman microspectroscopy provides a phase-specific, quantitative readout that links to ultrafast laser writing conditions and device performance in diamond.
+Across surface pads and wire electrodes, depletion of the 1332 cm-1 sp3 line predicts conductivity more reliably than absolute sp2 and PL intensities.
+Some of these deficiencies could be solved via spectral pre-processing, which has the drawback of adding complexity, time, and additional sources of error [chen2018fast].
+Therefore, a self-referenced sp3 metric, benchmarked to the local pristine background, is therefore a robust in situ indicator, while the sp2/sp3 ratio remains useful at moderate doses since it is less sensitive to changes in experimental conditions.
+Thus, Raman microspectroscopy can be integrated into the fabrication system to provide real-time, in-situ feedback of the device fabrication.
+In this integrated application, rather than generating the maps shown in this work, the excitation laser could be spatially offset with a lag with the fabrication laser and sample scattering could be collected with sub-second integration times.
+Additionally, as shown in Supporting Note 2, the excitation beam would contribute to removing surface debris.
+As part of this work, linear hyperspectral unmixing has been introduced as a full-spectrum, label-free alternative for post-fabrication mapping and indicator validation.
+In the case of diamond surface graphitization, unmixing performs analogously with band-based metrics, demonstrating its potential for samples with less defined structure-specific bands.
+The methodology presented in this work is not specific to diamond structure indicator and could therefore be adapted to other hosts and functionalities (e.g., fabrication in glasses and wide-bandgap crystals [fernandez2014role, Li2008, Little2010], laser-induced graphene on polymers [Wang2025], and embedded heaters or electrodes in SiC, sapphire, silica, and ceramics [Choi2016, Naseri2022, Zhang2025]), providing a practical route toward specification‐driven fs‐laser microfabrication.
+4 Materials and Methods
+4.1 Electrode fabrication
 A single-crystal, optical-grade diamond grown by chemical-vapor deposition (Element Six; nominal N concentration ppm) served as the substrate for this study.
-Surface graphitic pads and wires were inscribed with a fs laser operating at 520 nm ( fs, kHz) with various laser scan speed.
+Surface graphitic pads and wires were inscribed with a fs laser operating at 520 nm ( fs, kHz,) and a pulse energy of 30 nJ, using various laser scan speed.
 Wave-front aberrations were corrected by a liquid-crystal spatial light modulator, and the beam was focused through an objective lens with numerical aperture (ZEISS Plan-Neofluar, 0.5NA, 20x).
@@
 Pad electrodes were produced by stacking parallel wires with a constant 0.5 m pitch, maintained for all scan speeds.
-2 Electrical conductivity measurement
+4.2 Electrical conductivity measurement
 The resistance of laser-written pads and wires was determined from two-probe current–voltage (I–V) measurements using a low-noise transimpedance amplifier (DLPCA-200, FEMTO) together with a real-time voltage source and data-acquisition controller (ADwin-Gold II, Jäger).
@@
 The ADwin generated quasi-static voltage sweeps within the ohmic window (typically mV) and synchronously digitized the DLPCA-200 output.
-3 Brightfield imaging
+4.3 Brightfield imaging
 Optical micrographs were captured with an integrated transmission optical microscope.
 A red-LED array situated under the sample serves as illumination source and the images are captured on a CCD camera (EC650, Prosilica Inc.).
-4 Confocal Imaging
+4.4 Confocal Imaging
 Our fabrication setup has an integrated custom confocal microscope that can be simultaneously utilized for photoluminescence and Raman spectral collection.
-Using a 532 nm continuous-wave excitation laser (Cobolt Samba 150mW) and a fast steering mirror (FSM-300, Newport) relayed via a 4f into the back aperture of the system’s objective, 2D PL and RS maps can be collected.
-Notably, in this study, the 2D maps were collected post-fabrication utilising a higher numerical aperture objective (Olympus PlanApo 0.95NA, 60x).
-Both photoluminescence and Raman spectra were collected simulataneously employing 30 mW of laser excitation.
+Using a 532 nm continuous-wave excitation laser (Cobolt Samba 150mW) and a fast steering mirror (FSM-300, Newport) relayed via a 4f into the back aperture of the system’s objective, 2D photoluminescence and Raman maps can be collected.
+Notably, in this study, the 2D maps were collected post-fabrication utilizing a higher numerical aperture objective (Olympus PlanApo 0.95NA, 60x).
+Both photoluminescence and Raman spectra were collected simultaneously employing 30 mW of laser excitation.
 The samples were scanned with a 500nm pixel-to-pixel lateral displacement and a 1 second exposure time.
@@
 For the Raman collection path, scattered light was dispersed by a 1200 lines mm-1 blazed spectrograph (SpectraPro HRS-1200) and detected on a deep-cooled CCD camera (PIXIS 100, Princeton Instruments).
-5 Raman spectra preprocessing
+4.5 Raman spectra preprocessing
 The hyperspectral Raman data were loaded as a three-dimensional array .
@@
 Fixed integration windows were used throughout (sp3 width cm-1; sp2 width cm-1); thus, the sp2/sp3 ratio is the ratio of integrated band intensities, not intensity per unit wavenumber.
-Full algorithmic details, parameter choices, and examples are provided in the SI (see Supplementary Note 3).
-6 Spectral Unmixing
+Full algorithmic details, parameter choices, and examples are provided in the SI (see Supporting Note 3).
+4.6 Spectral Unmixing
 In this work we adopt the widely used linear mixing model (LMM), in which each spectrum is expressed as a linear combination of endmember spectra :
@@
 Endmember identification.
-To identify the endmember spectra from the data, we employed two well-established blind unmixing algorithms (a)Vertex Component Analysis (VCA) [41], a fast geometrical method that projects the dataset along random directions to iteratively identify extreme points of the spectral simplex, which serve as candidate endmembers; and (2) N-FINDR [42], an algorithm that searches for the set of endmembers maximising the volume of the simplex formed by candidate spectra, under the assumption that pure pixels are present in the dataset.
+To identify the endmember spectra from the data, we employed two well-established blind unmixing algorithms (a)Vertex Component Analysis (VCA) [nascimento2005vertex], a fast geometrical method that projects the dataset along random directions to iteratively identify extreme points of the spectral simplex, which serve as candidate endmembers; and (2) N-FINDR [winter1999n], an algorithm that searches for the set of endmembers maximising the volume of the simplex formed by candidate spectra, under the assumption that pure pixels are present in the dataset.
 Both algorithms require specifying the number of endmembers in advance.
@@
 Abundance estimation.
-Once endmembers are identified, abundances can computed by solving the linear inverse problem under different constraints: (a) Fully Constrained Least Squares (FCLS) [43]: enforces both non-negativity (), and the sum-to-one constraint (), ensuring abundances can be interpreted as fractions; and (b) Non-Negative Least Squares (NNLS) [44]: relaxes the sum-to-one condition but maintains non-negativity, which can be useful when baseline variations are present.
+Once endmembers are identified, abundances can computed by solving the linear inverse problem under different constraints: (a) Fully Constrained Least Squares (FCLS) [heinz2001fully]: enforces both non-negativity (), and the sum-to-one constraint (), ensuring abundances can be interpreted as fractions; and (b) Non-Negative Least Squares (NNLS) [lawson1995solving]: relaxes the sum-to-one condition but maintains non-negativity, which can be useful when baseline variations are present.
 As shown in Figure 3, the abundance maps at 50 m/s, the pristine diamond region does not show an abundance of exactly 0 for the graphite endmember and 1 for the diamond endmember.
-This is expected since the endmembers are extracted across all wire maps, focus and intensity differences between datasets also shift the estimated simplex (see Supplementary Note 8).
+This is expected since the endmembers are extracted across all wire maps, focus and intensity differences between datasets also shift the estimated simplex (see Supporting Note 8).
 The result is a small residual assignment of sp2 abundance even in pristine regions.
@@
 That said, three effects can challenge strict linearity: (i) strong broadband fluorescence in highly damaged zones, (ii) self-absorption and re-emission in thick, strongly absorbing tracks, and (iii) spectral distortions from stress/strain gradients that shift or broaden peaks in a way not representable by a fixed endmember.
-In such cases, non-linear or bilinear models (e.g., polynomial post–nonlinear mixing, kernel-based unmixing, or autoencoder-based unmixing [45]) may yield incremental gains, at the cost of additional parameters and reduced interpretability.
+In such cases, non-linear or bilinear models (e.g., polynomial post–nonlinear mixing, kernel-based unmixing, or autoencoder-based unmixing [georgiev2024hyperspectral]) may yield incremental gains, at the cost of additional parameters and reduced interpretability.
 Implementation.
-We conducted the hyperspectral unmixing analysis with Python using the RamanSPy package [46].
-V Acknowledgments
+We conducted the hyperspectral unmixing analysis with Python using the RamanSPy package [georgiev2024ramanspy].
+5 Acknowledgments
 This work was supported by the UK Engineering and Physical Sciences Research Council (EP/W025256/1).
-VI Disclosures
+6 Disclosures
 The authors declare no conflicts of interest.
-VII Data availability
+7 Data availability
 The data that support the findings of this study are available from the corresponding authors, [X.C., A.F.G.], upon reasonable request.
-Supplementary Information
-Supplementary Note 1: Pad optical micrograph and Raman map
-As desciribed in the manuscript, we employed laser-written graphitic pads and wires to emulate two common classes of conductive elements.
-Supplementary Figure 5a shows an optical micrograph of pad electrodes with a width of 50 m and length of 200 m, written at scan speeds of 100 and 200 m s-1, respectively.
+Supporting Information
+Supporting Note 1: Pad optical micrograph and Raman map
+As described in the manuscript, we employed laser-written graphitic pads and wires to emulate two common classes of conductive elements.
+Supporting Figure 1a shows an optical micrograph of pad electrodes with a width of 50 m and length of 200 m, written at scan speeds of 100 and 200 m s-1, respectively.
 The 100 m s-1 pad exhibits a resistance of , whereas the 200 m s-1 pad shows , i.e., a difference of more than an order of magnitude.
 Under optical illumination, the laser-written tracks appear dark relative to the surrounding single-crystal diamond.
-This contrast arises because converting transparent, wide-band-gap sp3 diamond to disordered sp2 carbon increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [24, 25].
-By contrast with the graphitic wires discussed in the main text, whose resistance difference is only weakly discernible in optical images, the two pads are optically indistinguishable.
-Supplementary Figures 5b and c present the corresponding photoluminescence (PL) and Raman (sp3 and sp2) maps.
+This contrast arises because converting transparent, wide-band-gap sp3 diamond to disordered sp2 carbon increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [prelas1997handbook, Wang2000].
+Supporting Figures 1b and c present the corresponding photoluminescence (PL) and Raman (sp3 and sp2) maps.
 In both pads, the written regions appear bright in PL immediately after fabrication.
 This behavior differs from the wires, which appear dark in PL due to laser-induced non-radiative recombination pathways.
-The bright pad emission consists of locally activated NV0 and NV- centres in this Type Ib diamond (high nitrogen content), and other defects centres that producing a broad band that photo-bleaches under green (532 nm) excitation.
+The bright pad emission consists of locally activated NV0 and NV- centers in this Type Ib diamond (high nitrogen content), and other defects centers that producing a broad band that photo-bleaches under green (532 nm) excitation.
 This highlights that monitoring the full-window PL signal is not an ideal metric during fabrication, as relevant phase information is obscured by unrelated emissions and does not directly reflect device performance.
 In contrast, hyperspectral Raman mapping provides more specific structural information.
-By integrating the sp3 diamond window (1325-1340cm-1) [26] and the sp2 G-band window (1575-1610cm-1) [24], the sp3 map shows the laser-written regions as decreased peak intensity relative to pristine diamond, with stronger depletion for the lower scan speed.
-The sp2 map exhibits, while exhibiting the expected behavior (i.e., increased peak intensity in the written region), has lower overall contrast.
+By integrating the sp3 diamond window (1325-1340cm-1) [Kononenko2008] and the sp2 G-band window (1575-1610cm-1) [prelas1997handbook], the sp3 map shows the laser-written regions as decreased peak intensity relative to pristine diamond, with stronger depletion for the lower scan speed.
+The sp2 map exhibits the expected behavior (i.e., increased peak intensity in the written region) but has lower overall contrast.
 Both the sp3 and sp2 Raman maps exhibit horizontal line-like features, which arise from the writing process.
 Such features are more pronounced in the sp2 than in the sp3 maps.
-Details explanation will be discussed in Supplementary Note 2.
-Supplementary Figure 6 shows an optical micrograph of a pad acquired after PL mapping.
-Scanning a 20 mW, 532 nm beam around the electrode removes the initial femtosecond-laser-induced graphite debris [25], revealing a clean structure (dashed green square) compared to the post-writing state (dashed red square).
+Details explanation will be discussed in Supporting Note 2.
+Supporting Figure 2 shows an optical micrograph of a pad acquired after PL mapping.
+Scanning a 20 mW, 532 nm beam around the electrode removes the initial femtosecond-laser-induced graphite debris [Wang2000], revealing a clean structure (dashed green square) compared to the post-writing state (dashed red square).
 Thus, maintaining the 532 nm probe during Raman acquisition could serve the dual role of (a) enabling spectroscopy monitoring, and (b) clearing graphite debris.
-With this, debris would not mask the written features during in situ monitoring, and no additional cleaning step is required.
-Supplementary Note 2: Origin of contrast in hyperspectral maps
-As shown in Supplementary Figure 7, both the sp3 and sp2 Raman maps exhibit horizontal line-like features, which are more pronounced in the sp2 map than in the sp3 map.
+Supporting Note 2: Origin of contrast in hyperspectral maps
+As shown in Supporting Figure 3, both the sp3 and sp2 Raman maps exhibit horizontal line-like features, which are more pronounced in the sp2 map than in the sp3 map.
 The origin of these contrasts differs for the two cases.
@@
 Here, the increased brightness is not related to enhanced graphitization but instead originates from broadband background emission.
-To confirm this, Supplementary Figure 7 presents a Raman map in the 590–600 nm spectral range (the same hyperspectral map of Figure 1c in the main text but different spectral window), excluding the sp3 and sp2 bands.
+To confirm this, Supporting Figure 3 presents a Raman map in the 590–600 nm spectral range (the same hyperspectral map of Figure 1c in the main text but different spectral window), excluding the sp3 and sp2 bands.
 Compared to the laser-written wire and the pure diamond background, the surrounding damaged shell appears bright, confirming that the observed contrast does not originate from either the sp3 or sp2 Raman signals.
-Spectra from three representative pixels—assigned to the graphite core, the damaged shell, and the pure diamond background—are extracted in Supplementary Figure 7a and shown in Supplementary Figure 7b–d.
-Each spectrum is self-normalized for comparison.
+Spectra from three representative pixels—assigned to the graphite core, the damaged shell, and the pure diamond background—are extracted in Supporting Figure 3a and shown in Supporting Figure 3b–d.
+Each spectrum is normalized for comparison.
 The graphite region exhibits depleted sp3, a clear D band, and sp2 features, while the pure diamond spectrum is dominated by the sp3 line.
@@
 Instead, a broad emission is present, which contributes to the increased brightness observed in both the photoluminescence and hyperspectral maps.
-This emission is attributed to defect activation in the damaged shell surrounding the wire core [26], leading to an overall increase in count rates.
-Supplementary Note 3: Pixel identification, integration, and normalization
-The hyperspectral Raman data was loaded as a four-dimensional array , where and represent the spatial coordinates, denotes the focal depth, and represents the wavelength.
-For all analyses, we selected the -plane closest to µm, effectively reducing the data to a three-dimensional array .
+This emission is attributed to defect activation in the damaged shell surrounding the wire core [Kononenko2008].
+Supporting Note 3: Pixel identification, integration, and normalization
+The hyperspectral Raman data was loaded as a three-dimensional array , where and represent the spatial coordinates, and represents the wavelength.
 To compensate for experimental variations including focusing, and excitation laser power fluctuations between different datasets, we implemented a sample-specific normalization procedure based on pristine diamond regions (shown as green pixels in Fig.
-8).
+4).
 For the sp3 diamond peak analysis, we defined a wavelength window cm-1. The integrated intensity for each pixel was calculated as:
@@
 Subsequently, we extracted and performed integrated intensity calculations for both the sp3 (diamond) and sp2 (graphitic carbon) spectral windows for those with dark sp3 intensity (red pixels in Fig.
-8).
+4).
 For each pixel, we computed:
@@
 The sp2/sp3 ratio should therefore be interpreted as the ratio of integrated intensities over their respective characteristic spectral features, rather than intensity per unit wavelength.
-Supplementary Note 4: Quantitative PL analysis
-To evaluate whether broadband photoluminescence (PL) intensity can serve as an in situ process indicator—analogous to the sp3, sp2 signals and their ratio, we extracted the PL intensity from the written pixels.
-The results are presented in Supplementary Figure 9.
+Supporting Note 4: Quantitative PL analysis
+To evaluate whether broadband photoluminescence (PL) intensity can serve as an in situ process indicator, we extracted the PL intensity from the written pixels.
+The results are presented in Supporting Figure 5.
 This figure shows a correlation between the resistance and the normalized PL intensity of the written wire electrodes, which suggests that broadband PL intensity can be leveraged to assess performance of the written electrodes.
-However, given the broad spectral collection window of these PL maps (550-850 nm), they can suffer from significant fluctuations due to factors unrelated to the phenomenon of interest, in this case, graphitization.
-For example, in Type Ib diamond, however, intrinsic defects generate a non-negligible background in PL maps; consequently, the brightest and darkest PL pixels do not directly correspond to “pure diamond” or “graphitized” regions.
-Supplementary Figure 9, in this particular case, since the sp3 component dominates the broadband PL for the written region, the normalized PL follows the sp3 trend: PL increases with resistance, consistent with a larger fraction of diamond-like bonding.
-Nevertheless, the broadband PL also includes emission from defect clusters, leading to higher uncertainty (i.e.
-measurement variance) than for the sp3 metric.
-In summary, broadband PL can serve as a quantitative indicator during laser fabrication, but it is less specific and less accurate than the sp3 intensity.
-Accordingly, spectrally isolating the sp3 (and/or sp2) Raman windows—rather than relying on fully broadband PL—provides the most reliable and clean in situ process feedback.
-Supplementary Note 5: sp3 distribution
-The speed-dependent probability-density histograms of the normalised sp3 Raman intensity, computed from pixels identified as graphitised in the hyperspectral maps, are shown for pad and wire electrodes in Supplementary Figure 10 and Supplementary Figure 11, respectively.
-For each scan speed, 1500 pixels are analysed for pads and 60 for wires.
-Intensities are referenced to pristine diamond, so higher normalised sp3 corresponds to weaker graphitization.
+Supporting Figure 5 also compares the PL with the sp3 intensity metric proposed in the main manuscript.
+This comparison showcases how the normalized PL follows the sp3 trend: PL increases with resistance, consistent with a larger fraction of diamond-like bonding.
+Yet, despite having similar correlation with the resistance of the electrodes, the PL normalized intensity presents significantly larger variance in the measurements.
+This is due to the broad spectral collection window of these PL maps (550-850 nm), which can include emission from phenomena not directly related to graphitization (i.e., in Type Ib diamond intrinsic defects generate a non-negligible background in PL maps).
+Thus, even if broadband PL can serve as a quantitative indicator during laser fabrication, in practice it is less specific and less accurate than the sp3 intensity, which should be favored for in situ process monitoring.
+Supporting Note 5: sp3 distribution
+The speed-dependent probability-density histograms of the normalized sp3 Raman intensity, computed from pixels identified as graphitized in the hyperspectral maps, are shown for pad and wire electrodes in Supporting Figure 6 and Supporting Figure 7, respectively.
+For each scan speed, 1500 pixels are analyzed for pads and 60 for wires.
+Intensities are referenced to pristine diamond, so higher normalized sp3 corresponds to weaker graphitization.
 In pad electrodes, increasing scan speed shifts the distributions toward higher values and simultaneously broadens them, consistent with reduced energy deposition per unit length and a less uniform sp3-to-sp2 transformation.
 By contrast, wire electrodes exhibit narrower distributions across all scan speeds because their geometry prevents overwriting.
-Nonetheless, they display the same overall trend as pads: increasing scan speed shifts the distributions to higher normalised sp3 values, reflecting a lower degree of graphitization in the laser-written structures.
-Supplementary Note 6: sp3 Raman analysis
+Nonetheless, they display the same overall trend as pads: increasing scan speed shifts the distributions to higher normalized sp3 values, reflecting a lower degree of graphitization in the laser-written structures.
+Supporting Note 6: sp3 Raman analysis
 To complement the two-probe transport data, we quantified how the sp3 Raman response varies with laser scan speed.
 For each speed, 60 pixels within the graphitized region of a laser-written wire were averaged and the sp3 peak was fitted with a Lorentzian.
-The wire geometry was chosen over the pad because it provides a cleaner, more localized transformation without overlapping features.
 Averaged spectra for scan speeds from 10 to 1270 m s-1 were analyzed; the resulting peak center and full width at half maximum (FWHM) are plotted in Fig.
@@
 A pristine-diamond reference, obtained by averaging 100 pixels from an unexposed area, is shown at 3000m s-1 for visual comparison only.
-Across the series, the sp3 mode lies slightly above 1332cm-1, consistent with tensile stress in the transformed graphite layer and the concomitant in-plane compressive stress imparted to the adjacent diamond[25], which produces a blue shift of the diamond Raman mode [31, 32].
+Across the series, the sp3 mode lies slightly above 1332cm-1, consistent with tensile stress in the transformed graphite layer and the concomitant in-plane compressive stress imparted to the adjacent diamond[Wang2000], which produces a blue shift of the diamond Raman mode [Grimsditch1975, Zaitsev2001].
 The FWHM decreases systematically with increasing scan speed.
 Slower scans deposit more energy per unit length, driving a higher degree of graphitization; accordingly, the sp3 line departs from an ideal Lorentzian, broadening and sometimes becoming asymmetric.
-This behavior reflects increasing lattice disorder and strain distributions (inhomogeneous broadening), reduced phonon lifetimes (homogeneous broadening), and phonon-confinement effects; if the material becomes sufficiently conducting, Fano interference with an electronic continuum can further skew the lineshape [33, 47].
+This behavior reflects increasing lattice disorder and strain distributions (inhomogeneous broadening), reduced phonon lifetimes (homogeneous broadening), and phonon-confinement effects; if the material becomes sufficiently conducting, Fano interference with an electronic continuum can further skew the lineshape [Orwa2000, Orwa2009].
 All averaged spectra and their Lorentzian fits are provided in Fig.
-13.
-Supplementary Note 7: Electrical performance for both pad and wire
+9.
+Supporting Note 7: Electrical performance for both pad and wire
 Electrical measurements confirm that femtosecond-laser writing produces ohmic graphitic pathways whose conductance scales with scan speed and tracks the depletion of sp3 Raman intensity in the written region.
 For wire electrodes, the I-V characteristics are strictly linear from -0.5 to +0.5 V, and linear fits yield resistances that vary over several orders of magnitude as the scan speed increases from 10 to 1000 m s-1 (Fig.
-14a-l).
+10a-l).
 At the slowest scans (10-50 m s-1) the wires exhibit resistances in the tens of k, approaching the series resistance floor of the test board contact pads, whereas progressively faster scans—depositing less energy per unit length—produce markedly higher resistances reaching the M-G range.
@@
 I-V curves acquired from -0.1 to +0.1 V remain perfectly linear, with fitted resistances rising from a few hundred at 100 m s-1 to 50k at 500 m s-1 (Fig.
-15a-e).
+11a-e).
 The comparison between wires and pads highlights the expected geometric scaling and the writing process: wider, thicker pads support significantly higher conductance than narrow wires under otherwise identical conditions, and the zig-zag writing strategy introduces controlled overwriting that further enhances graphitization.
@@
 This direct Raman-electrical correlation substantiates that efficient graphitization—conversion from sp3 to sp2 bonding—governs charge transport in the written structures.
-Supplementary Note 8: Hyperspectral unmixing
-We performed hyperspectral unmixing with a linear mixing model using VCA for endmember extraction and FCLS for abundance estimation.
-As shown in Supplementary Figure 16, the unmixing results consistently show progressive loss of the sp3-like component with decreasing scan speed, while the sp2-like component increases and eventually saturates.
+Supporting Note 8: Hyperspectral unmixing
+We performed hyperspectral unmixing with a linear mixing model using vertex component analysis (VCA) [nascimento2005vertex] for endmember extraction and fully-constrained least squares (FCLS) for abundance estimation.
+As shown in Supporting Figure 12, the unmixing results consistently show progressive loss of the diamond-associated component with decreasing scan speed, while the carbon-associated component increases and eventually saturates.
 This reinforces the main conclusion that sp3 depletion provides a reliable, monotonic indicator of graphitization.
@@
 However, it is sensitive to the presence of noise or species not considered in the analysis.
-As shown in Supplementary Figure 16a, the presence of the broadband emission damage shell when fabricating at intermediate speeds is mislabeled as being a mixture of diamond and graphite phases.
-Supplementary Figure 17 presents the endmembers extracted using different unmixing techniques and spectral regions of interest, exhibiting the variance that these endmembers can have.
-This behavior is expected: VCA and N-FINDR are geometric algorithms that approximate the data as lying within a convex simplex.
-They perform best when pixels lie close to the true extremes, but in practice the confocal voxel contains mixtures, variable SNR, NV0 fluorescence, and stress-induced peak shifts.
-The algorithms therefore identify different “extreme” representatives of the sp3-like and sp2-like responses, leading to small method-dependent variations that do not alter the physical interpretation.
+As shown in Supporting Figure 12a, the presence of the broadband emission damage shell when fabricating at intermediate speeds is mislabeled as being a mixture of diamond and graphite phases, since in the analysis only two endmembers are considered.
+Supporting Figure 13 presents the endmembers extracted using different unmixing techniques and spectral regions of interest (ROI), exhibiting the variance that these endmembers can have.
+This behavior is expected: VCA and N-FINDR [winter1999n] are geometric algorithms that approximate the data as lying within a convex simplex.
+They perform best when pixels lie close to the true extremes, but in practice the confocal voxel contains mixtures, variable SNR, NV0 fluorescence, and potentially stress-induced peak shifts.
+The algorithms therefore identify different “extreme” representatives of the diamond-like and graphite-like responses, leading to small method-dependent variations that do not alter the physical interpretation.
 Indeed, the endmembers preserve the spectral features of their associated spectra, that is, the sp3 peak and NV0 band for the endmember associated to diamond, and the G band and, to lesser extent, the D band for the one associated to graphite.
-Notably, for the ROI including the sp3 peak, the large dynamic range of this feature amplifies small fluctuations which can distort the convex geometry and bias endmember selection.
-Excluding this line (ROI 1345–1800 cm-1) reduces leverage and produces noisier but qualitatively similar maps.
-It is important to distinguish geometric unmixing from PCA.
+Notably, for the spectral ROI including the sp3 peak, the large dynamic range of this feature amplifies small fluctuations which can distort the convex geometry and bias endmember selection.
+Excluding this line (spectral ROI: 1345–1800 cm-1) reduces leverage and produces noisier but qualitatively similar maps.
+Supporting Figure 13 also provides the first two principal components.
+It is important to distinguish geometric unmixing from principal component analysis (PCA) [abdi2010principal].
 VCA and N-FINDR seek spectra at the boundaries of the data cloud that can be interpreted as approximate endmembers, whereas PCA decomposes variance into orthogonal components that need not resemble real spectra and can take negative values.